### v1

# OWT: A Foundational Organ-Wise Tokenization Framework for Medical Imaging

Abstract
Recent advances in representation learning often rely on holistic, black-box embeddings that entangle multiple semantic components, limiting interpretability and generalization. These issues are especially critical in medical imaging. To address these limitations, we propose an Organ-Wise Tokenization (OWT) framework with a Token Group-based Reconstruction (TGR) training paradigm. Unlike conventional approaches that produce holistic features, OWT explicitly disentangles an image into separable token groups, each corresponding to a distinct organ or semantic entity. Our design ensures each token group encapsulates organ-specific information, boosting interpretability, generalization, and efficiency while allowing fine-grained control in downstream tasks. Experiments on CT and MRI datasets demonstrate the effectiveness of OWT in not only achieving strong image reconstruction and segmentation performance, but also enabling novel semantic-level generation and retrieval applications that are out of reach for standard holistic embedding methods. These findings underscore the potential of OWT as a foundational framework for semantically disentangled representation learning, offering broad scalability and applicability to real-world medical imaging scenarios and beyond.  [^1] Code is available at https://github.com/SifanSong/OWT
1 Introduction
Recent years have witnessed significant advances in representation learning of visual data for both natural and medical images. The typical representation learning, driven by self-supervised and task-guided training frameworks, is classified as holistic embedding-based representation learning since it learns embeddings that capture all information for an entire image [54, 22, 58] (Fig.1-a). Similar trends appear in the medical imaging domain, where holistic representations have also been extensively explored for tasks like disease diagnosis[49], organ segmentation [46, 5, 25], and anatomical structure analysis [41, 59]. Although the holistic embeddings leveraged by these studies have been proven effective, they remain globally entangled and mixed in a black-box latent space, combining information about multiple organs and background. This fusion decreases efficiency and poses significant challenges for organ-focused analyses, disease diagnosis, or radiological interpretation, where fine-grained interpretability and a semantically targeted focus are critically important [48, 6, 58].
Figure 1: Objective comparison of (a) holistic embedding-based representation learning, (b) disentangled representation learning, and (c) our OWT framework. Unlike prior methods, our foundational tokenization framework enables the extraction of disentangled token groups containing semantically meaningful organ-specific representations. These token groups can be leveraged separately or in combination for downstream analyses, such as semantic-level reconstruction, segmentation, and retrieval.
In parallel, disentangled representation learning (DRL) has been studied to decompose data into semantically meaningful factors of variation, with each factor ideally corresponding to an independent aspect of the underlying data [58] (Fig.1-b). By isolating latent factors (e.g., shape, color, orientation) in natural images [4, 12, 61, 64] or modality attributes in medical images [44, 8], DRL methods aim to improve explainability and enable controllable feature manipulation. However, most conventional DRL approaches rely on purely data-driven or perturbation-driven strategies, requiring extensive modeling assumptions or specialized regularizers to ensure each factor corresponds to a semantically coherent concept [38, 58]. Consequently, existing DRL approaches face notable challenges, including limited alignment with real-world semantics, weak spatial correspondence, and poor scalability efficiency.
Despite these successes, existing holistic embedding–based strategies and DRL approaches still have notable gaps, particularly in medical imaging, where high-level interpretability and organ-specific analysis are crucial (e.g., anatomical analysis and surgical planning). Specifically, existing methods suffer from (a) a lack of interpretability and generalizability, as multiple semantic cues are often merged without alignment to real-world concepts, making it difficult to isolate relevant structures from confounding factors (e.g., background variations); and (b) limited efficiency and controllability, since these approaches process entire or masked representations rather than focusing only on organ-level features, restricting the ability to selectively manipulate or analyze distinct semantic factors.
These limitations motivate us to revisit the design of representation learning, aiming for explicit organ- (or part-) level tokenization to tackle them. In this paper, we propose a foundational Organ-Wise Tokenization (OWT) framework and a novel Token Group-based Reconstruction (TGR) training paradigm. Specifically, in addition to a standard encoder-decoder architecture, such as Masked Autoencoders (MAE) [27], our OWT framework incorporates two key novel modules, Organ Collector and Adaptive Holistic Embedding Restorer (AHER), inspired by prior structures for extracting semantic features [47, 18]. The Organ Collector integrates holistic embeddings into semantically meaningful token groups, while AHER reconstructs holistic representations from any number of token groups. Together, these modules enable a structured transition from holistic $\rightarrow$ semantic $\rightarrow$ holistic embeddings, facilitating both organ-specific processing and reconstruction. Meanwhile, our proposed TGR training paradigm ensures that the extracted semantic embeddings are fully functional as Semantically Disentangled Token Groups (SDTGs) by explicitly correlating each token group to its corresponding organ. In this case, each SDTG encapsulates only organ-specific information and is capable of being utilized independently or in any combination with other SDTGs as practically needed (Fig. 1-c).
Therefore, our design offers key advantages by addressing the limitations of existing approaches: (i) improved interpretability and generalizability, as each token group is explicitly associated with a specific organ or semantic content, enabling the isolation of target structures while minimizing confounding factors; and (ii) enhanced efficiency and controllability, as OWT processes only the relevant token groups for a given task, reducing computational load and allowing for semantic-based, fine-grained manipulation and adaptation without introducing unrelated features.
Figure 2: Overall architecture of Organ-Wise Tokenization (OWT). The input 2D or 3D images are first encoded into holistic embeddings ($X_{H}$). Then, the holistic tokens are disentangled and embedded in an organ-wise manner using the Organ Collector, forming Semantically Disentangled Token Groups (SDTGs, $X_{G}$). The SDTGs with solid lines represent randomly retained token groups ($\tilde{X}_{G}$). The Token Group Encoder further processes these retained groups, capturing both inter- and intra-organ relationships, and transforming them into $\tilde{X}^{\prime}_{G}$. Finally, an Adaptive Holistic Embedding Restorer and a decoder sequentially restores the holistic representation ($X^{\prime}_{H}$) and reconstructs the final output ($\hat{I}_{T}$), enabling semantic-level reconstruction.
Figure 3: Comparison of the overall process of holistic embedding-based and token group-based reconstruction.
We evaluate the performance of the organ-wise embeddings (i.e., SDTGs) extracted by OWT across multiple downstream tasks (such as image reconstruction, segmentation and retrieval). Across multiple modalities (CT, two MRI sequences, and a facial dataset), our method achieves better and more robust results compared to baselines (i.e., MAE [27] and dedicated semantic segmentation networks [63, 42, 9, 5, 51]) and is able to accomplish tasks that these traditional methods cannot, such as semantic-level generation, retrieval and embedding combination. This proves the concept of our foundational semantically disentangled tokenization framework, enabling us to bring more potential solutions for real-world applications.
In summary, our contributions are listed as follows:
  • We propose a foundational framework and a novel training paradigm, OWT and TGR, which encourage the network to encapsulate organ-level features into semantically disentangled token groups (SDTGs) rather than entangled embeddings with a black-box nature.
  • The proposed SDTGs are highly condensed with rich information on each target organ, which increases the interpretability and generalization of latent embeddings.
  • The token groups can be used separately or in combination and are processed with improved efficiency and controllability for downstream tasks and further analysis.
  • Experiments demonstrate that our method offers significant scalability. It can adaptively integrate semantic information based on various requirements and incorporate other modalities (e.g., text encoding), highlighting its potential for broader applications beyond medical imaging.
2 Related Work
2.1 Holistic Embedding-based Representation Learning
Holistic embedding-based approaches map entire 2D/3D inputs into a single latent representation, aiming to capture global image features [54, 22, 58]. These methods typically follow either self-supervised or task-guided strategies. In self-supervised learning, contrastive techniques like SimCLR [10], SimSiam [11], and BYOL [24] enforce consistency between augmented views, while predictive approaches such as CPC [43] and RotNet [23] reconstruct missing or permuted portions. Recently, masked image modeling (e.g., MAE [27] and BEiT [3]) has emerged as an effective approach by requiring models to infer masked regions. Meanwhile, task-guided frameworks focus on extracting features that directly benefit a target task, typically resulting in holistic embeddings. Many classic neural networks, such as ResNet [26], DenseNet [29], and Vision Transformer [16], and their variants incorporate classification, regression, or segmentation objectives.
In medical imaging, holistic representations have been widely used for self-supervised feature learning in CT [13, 32], MRI [7, 33], and X-ray [52, 60]. Additionally, classical supervised pipelines (e.g., U-Net [46] and its variants [63, 9, 5, 25]) rely on a global bottleneck to embed entire images or volumes, demonstrating strong performance in organ segmentation and disease detection. Although such approaches achieve notable success, they often encode multiple semantic factors, such as multiple organs and the surrounding background, into an entangled representation with limited interpretability and fine-grained control [48, 6, 58].
2.2 Disentangled Representation Learning
In contrast to holistic embedding-based approaches, various disentangled representation learning (DRL) studies have pursued a dimension-wise strategy for two main objectives. First, to derive more robust holistic embeddings with improved generalizability [14, 1, 15, 56, 20, 50, 58], and second, to explicitly decompose images into latent factors, each corresponding to distinct and semantically meaningful attributes, in order to improve explainability [4, 12, 61, 64].
In medical imaging, the motivations for DRL are especially strong: improved interpretability, controllability, and robustness to confounding factors are crucial in clinical settings [37, 55]. When transitioning DRL to medical imaging, researchers seek to decouple anatomical properties (e.g., organ size or shape) [57, 30], modality characteristics (e.g., CT and MRI sequences) [44, 8], and pathological factors (e.g., tumor presence or severity) [36, 28], aiming to isolate distinct clinical features.
Despite notable progress, data- and perturbation-driven DRL approaches can be sensitive to training objectives, architecture choices, and assumptions about independent factors [38, 56, 58]. In existing studies, disentangled dimensions may not align precisely with real anatomical or pathological concepts, and spatial correspondence often remains weak when images are globally encoded. Consequently, there has been a growing emphasis on methods that combine disentanglement principles with controlled grouping or tokenization schemes to encapsulate semantically meaningful subsets of features.
3 Method
In this paper, we present a novel OWT framework and a Token Group-based Reconstruction (TGR) training paradigm, which embeds each organ into its corresponding semantically disentangled token group (SDTG). In this case, these organ-specific embeddings can be utilized independently or in combination for downstream tasks.
3.1 Overall OWT Framework
As illustrated in Fig. 2, our proposed OWT framework comprises five components: an encoder, Organ Collector, Token Group Encoder, Adaptive Holistic Embedding Restorer (AHER), and a decoder. The encoder maps input images to holistic embeddings, while the Organ Collector and TGR enable the target-organ features to be extracted and stored in specific embeddings (i.e., SDTGs). In TGR, we randomly select a subset of the predefined token groups. The retained token groups are fed into the Token Group Encoder to integrate both inter-organ and intra-organ information. Next, AHER enables any number of retained token groups to be converted back into a holistic embedding. Finally, the decoder reconstructs an image of the same size as the input.
The encoder and decoder in OWT follow standard holistic representation learning architectures, so they can be Transformer-based or (fused) CNN-based, as in 2D/3D MAE [27, 19] or 2D/3D VQGAN [17, 19]. This design choice allows the framework to be applied to either 2D or 3D images with minimal modification. For a 3D input volume processed by a Transformer-based encoder, let $I\in\mathbb{R}^{D\times H\times W\times C}$. The encoder ($\mathbf{Enc}$) produces a holistic embedding $X_{H}=\mathbf{Enc}(I)\in\mathbb{R}^{(dhw)\times c_{e}}$, where $D$, $H$, $W$ and $C$ represent the depth, height, width, and channels of the volume, respectively, while $d$, $h$ and $w$ are the corresponding embedded dimensions. The $(dhw)$ represents the number of tokens processed by the Transformer blocks, with each token having an embedding dimension of $c_{e}$. This holistic representation is then fed into subsequent modules for more fine-grained organ-wise processing.
3.2 Organ Collector
To effectively extract semantically meaningful features from the holistic embedding $X_{H}$, we introduce a novel Organ Collector module. A key component of this process is the token-wise attention matrix $A_{t}$, which dynamically assigns importance to different spatial locations or tokens, ensuring that each token group captures the most relevant semantic features.
To achieve this, let $X_{H}\in\mathbb{R}^{(dhw)\times c_{e}}$ be reshaped to $\mathbb{R}^{c_{e}\times d\times h\times w}$ before being fed into the Organ Collector (Fig. 2). We then employ a token-wise attention mechanism, defined as:
$A_{t}=\text{Softmax}(\text{Flatten}(\alpha(X_{H})))\;\in\;\mathbb{R}^{N_{g}\times(hw)},$ (1)
$X_{G}=A_{t}\,\text{Flatten}(\gamma(X_{H}))^{T}\;\in\;\mathbb{R}^{N_{g}\times c_{e}},$ (2)
where $N_{g}$ denotes the number of semantic-based tokens, and both $\alpha$ and $\gamma$ are linear layers. Concretely, $\alpha$ projects the $(c_{e}\times d)$-dimensional input to $N_{g}$, then applies Flatten and Softmax to derive a token-wise attention matrix. This attention matrix is multiplied by $\gamma(X_{H})$ (also flattened) to obtain $X_{G}\in\mathbb{R}^{N_{g}\times c_{e}}$.
Abdomen1k Holistic-based Rec. 1 Token Group Rec.
L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$ L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$
MAE [27] 6.74e-04 0.0684 0.9688 - - -
MAE/${}_{\text{LPIPS}}$ 6.19e-04 0.0403 0.9722 - - -
OWT 4.16e-04 0.0390 0.9734 3.31e-04 0.0073 0.9795
OWT/${}_{\text{4slices}}$ 3.69e-04 0.0371 0.9755 2.68e-04 0.0069 0.9838
AutoPet Holistic-based Rec. 1 Token Group Rec.
L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$ L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$
MAE [27] 2.86e-04 0.0452 0.9799 - - -
MAE/${}_{\text{LPIPS}}$ 2.77e-04 0.0255 0.9843 - - -
OWT 1.49e-04 0.0232 0.9881 3.32e-04 0.0071 0.9778
OWT/${}_{\text{4slices}}$ 1.13e-04 0.0179 0.9908 2.68e-04 0.0063 0.9827
Table 1: Performance of image reconstruction on CT datasets.
Abdomen1k Holistic-based Rec. 1 Token Group Rec.
L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$ L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$
MAE [27] 1.37e-03 0.0465 0.9686 - - -
MAE/${}_{\text{LPIPS}}$ 1.36e-03 0.0389 0.9690 - - -
OWT 7.65e-04 0.0256 0.9851 5.96e-04 0.0088 0.9762
OWT/${}_{\text{4slices}}$ 5.81e-04 0.0214 0.9885 5.04e-04 0.0081 0.9801
PreArtery Holistic-based Rec. 1 Token Group Rec.
L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$ L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$
MAE [27] 1.67e-03 0.0438 0.9701 - - -
MAE/${}_{\text{LPIPS}}$ 1.74e-03 0.0375 0.9672 - - -
OWT 8.66e-04 0.0248 0.9851 5.29e-04 0.0089 0.9759
OWT/${}_{\text{4slices}}$ 7.07e-04 0.0218 0.9876 4.47e-04 0.0082 0.9801
Table 2: Performance of image reconstruction on MRI datasets.
3.3 Token Group-based Reconstruction (TGR)
The TGR training paradigm provides direct supervision to token groups, ensuring their semantic disentanglement. By enforcing accurate organ-wise feature extraction and reconstruction through the token group-level loss $\mathcal{L}_{T}$ (Fig. 3-b), TGR preserves both spatial and textural consistency. Consequently, the extracted semantic tokens serve as semantically disentangled token groups (SDTGs). Through TGR, the Organ Collector adaptively extracts organ-specific information into its corresponding SDTG, assigning each group to a particular organ.
Definition of Token Group. Given a dataset $D$, we have its ground-truth annotations for $g$ organs. The remaining regions are treated as background and assigned to a single additional token group. Thus, we define a total of $(g+1)$ token groups. Let each group contain $f$ tokens, making the total number of tokens $N_{g}=(g+1)\times f$. For example, in Fig. 2, $X_{0}$, $X_{1}$, and $X_{2}$ represent the token groups that store liver, kidney, and background information, respectively.
Random Selection Procedure. To manage these token groups so that they are responsible for each organ during training, we adopt a two-stage random selection procedure. First, randomly determine the number of retained groups: Let $\Omega\sim\mathcal{U}(0,1)$ be a uniform random variable on $[0,1]$. With $(g+1)$ total Token Groups, we define $\tilde{g}=\bigl{\lfloor}(g+1)\cdot\Omega\bigr{\rfloor}$, resulting in an integer $\tilde{g}\in\{0,1,\dots,g\}$. Second, randomly select $\tilde{g}$ groups: From the full set of $(g+1)$ token groups, we then randomly pick exactly $\tilde{g}$ groups to retain and discard the rest. We set each group to contain $f$ tokens so that each separable and independent token group is in $\mathbb{R}^{{f}\times c_{e}}$. We concatenate the $\tilde{g}$ selected groups along the token dimension, producing the final retained tokens $\tilde{X}_{G}\in\mathbb{R}^{{\tilde{N}_{g}}\times c_{e}}$ (Fig 2), where the final retained token count is $\tilde{N}_{g}=\tilde{g}\times f$.
During TGR, for each input image, we remain only areas according to randomly retain token groups based on organ annotations (Fig. 3-b), treating the processed image as the target $I_{target}$ for the token group-based reconstruction loss $\mathcal{L}_{T}$. This explicit supervision encourages a precise organ-level focus, preserving each organ’s unique spatial and textural details in the retained token groups (i.e., SDTGs). Meanwhile, any randomly excluded token groups are discarded at this stage, ensuring that only the retained groups contribute to reconstruction and thereby enhancing semantic flexibility and computational efficiency. As illustrated in Fig. 3, TGR departs from conventional holistic embedding-based reconstruction methods that rely on the latent space for the entire image [2, 27, 19, 17] (Fig. 3-a). By selectively preserving certain token groups, our approach avoids fusing diverse organ features and does not require masked tokens, resulting in improved interpretability, efficiency, and controllability during both reconstruction and subsequent tasks.
After that, we employ a Token Group Encoder ($\mathbf{Enc}_{G}$) composed of six multi-head linear self-attention blocks [31]. Let $\tilde{X}^{\prime}_{G}=\mathbf{Enc}_{G}(\tilde{X}_{G})\in\mathbb{R}^{\tilde{N}_{g}\times c_{e}}$ represent the further contextual modeling that captures inter- and intra-organ relationships of the retained token groups.
3.4 Adaptive Holistic Embedding Restorer
Instead of relying on conventional masked tokens, our method discards unselected token groups in the random selection procedure, simplifying training and avoiding unnecessary overhead. After the last module, each input can produce a different number of retained token groups ($\tilde{X}^{\prime}_{G}$ in Fig. 2), resulting in a variable token count $\tilde{N}_{g}$. To convert these tokens back into a representation aligned with the original input size (i.e., the shape of the pre-Organ Collector embedding, $X_{H}$), we propose an Adaptive Holistic Embedding Restorer (AHER). This module dynamically accepts the variable token count $\tilde{N}_{g}$ and restores them to a consistent dimension suitable for subsequent reconstruction.
The AHER module is also implemented using a token-wise attention mechanism, defined as:
$A^{\prime}_{t}=\text{Softmax}(\phi(\tilde{X}^{\prime}_{G})^{T})\;\in\;\mathbb{R}^{(dhw)\times\tilde{N_{g}}}\;,$ (3)
$X^{\prime}_{H}=A^{\prime}_{t}\psi(\tilde{X}^{\prime}_{G})\;\in\;\mathbb{R}^{(dhw)\times c_{e}}\;,$ (4)
where $\phi(\cdot)$ and $\psi(\cdot)$ are two linear transformations acting on $\tilde{X}^{\prime}_{G}$, and $(dhw)$ is the flattened spatial dimension we seek to restore (i.e., the depth $\times$ height $\times$ width of the input volume). Specifically, a token-wise attention matrix $A^{\prime}_{t}$ is computed via softmax on the transposed $\phi(\tilde{X}^{\prime}_{G})$. Next, the features $\psi(\tilde{X}^{\prime}_{G})$ are re-aligned according to $A^{\prime}_{t}$, yielding a restored holistic embedding $X^{\prime}_{H}$ (Fig. 2)
Although $\tilde{N}_{g}$ varies per input, this operation preserves organ-specific features while ensuring $X^{\prime}_{H}$ remains a fixed $(dhw)\times c_{e}$ shape, thereby maintaining consistency with the decoder. Finally, we pass $X^{\prime}_{H}$ through a standard decoder [27] ($\mathbf{Dec}$) and unpatchify the output, reconstructing the image $\hat{I}_{T}=\mathbf{Dec}(X^{\prime}_{H})\in\mathbb{R}^{D\times H\times W\times C}$.
As mentioned in Sec. 3.3, our objective is to reduce the discrepancy between the target image $I_{target}$ and the reconstructed image $\hat{I}_{T}$. We combine an L2 term with the LPIPS perceptual loss [62]
$\mathcal{L}_{T}=\|I_{target}-\hat{I}_{T}\|_{2}^{2}\ +\mathcal{L}_{LPIPS}(I_{target},\hat{I}_{T})\;$ (5)
to preserves low-level fidelity while capturing high-level perceptual similarity, thereby guiding the network to generate more realistic texture and coherent organ regions.
4 Experiments
Figure 4: Samples of holistic-based and semantic-based medical image reconstruction from an MRI dataset.
4.1 Datasets and Experimental Setup
Datasets. We conduct experiments on four 3D medical datasets: two CT sets (Abdomen1k [40] and the CT volumes from AutoPet [21]) and two synthetic MRI sequences (delay and pre-artery) from RAOS dataset [39]. These multimodal and multi-sequence datasets validate the effectiveness and robustness of the OWT framework in medical imaging. To further validate that OWT can generalize to different semantic groupings beyond medical scenarios, we employ the CelebAMaskHQ dataset [34]. More details on the preprocessing steps are in Appendix 6.1.
Experimental Setup. All experiments are conducted on six Nvidia H100 GPUs. The hyperparameters, optimizer and scheduler settings follow MAE [27]. We compare OWT against a baseline MAE due to the similar encoder-decoder structure. In OWT, the default token number for each token group is 20. For processing 3D volumes (4 slices by default), we adopt a 3D version of MAE [19]. For fairness, we add LPIPS loss [62] to MAE to improve reconstruction quality. We also explore replacing the encoder-decoder pair with that of VQGAN [17]. Further details of architectures and hyperparameters are in Appendices 6.2 and 6.3, and ablation studies of default settings are in Appendix 8.
4.2 Results
Our OWT framework not only yields highly condensed representations containing rich information but also provides semantically meaningful and disentangled token groups. In this section, we design extensive experiments to evaluate OWT’s performance on several downstream tasks, including reconstruction, segmentation, generalization, and retrieval (Sec. 4.2.1). Notably, our findings indicate that, in contrast to conventional holistic-based methods, OWT can focus on semantic-level tasks and analyses, delivering superior interpretability and controllability. Next, we investigate the characteristics of these semantically disentangled token groups and explore their potential to incorporate other modalities (Sec. 4.2.2). We then demonstrate that the OWT framework can be employed in real-world applications (Sec. 4.2.3). Additionally, we provide ablation studies and explore more architectures of the OWT framework in Appendix 8.
4.2.1 Downstream Analysis
Image Reconstruction. We evaluate the reconstruction capability of OWT across four datasets (two CT and two MRI) [40, 21, 39]. Tables 1 and 2 show that OWT outperforms MAE on L2, LPIPS, and 3D SSIM metrics in holistic-based reconstruction tasks. Moreover, leveraging a patch-embedding module that tokenizes four slices for 3D volumes generates better results than the 2D OWT variant.
We further explore a semantic-based reconstruction scenario by retaining only a single organ-wise token group at a time, feeding it through AHER and the decoder to generate an organ-specific image. As reported in Tables 1 and 2, using a larger number of slices in CT and MRI volumes leads to more robust disentangled token groups, thereby improving semantic-based reconstruction quality. Fig. 4 illustrates MRI cases (rich in textural detail). Even with LPIPS added, MAE suffers from noticeable detail loss, whereas OWT (all token groups retained) produces sharper and clearer details. Additionally, we can freely combine the token groups to reconstruct only the content of interest.
In the zoom-in rows of Fig. 4, the liver/kidney reconstruction appears nearly identical whether we use only the liver/kidney token group or combine it with others, suggesting that OWT’s organ token groups are highly semantically independent. This property offers better interpretability as well as controllability, enabling reconstruction without unnecessary entanglement of unrelated organ features.
Figure 5: Semantic-based retrieval of CT/MRI images.
Methods CT MRI
Abdmen1k AutoPet Delay PreArtery
UNetplus [63] 88.38 89.29 84.11 84.42
AttU_Net [42] 89.34 89.73 84.54 85.04
TransUnet [9] 88.91 89.70 84.97 84.29
Swin-Unet [5] 86.56 88.06 83.07 82.41
CMUNet [51] 88.33 89.48 84.82 85.56
OWT 89.30 89.52 85.85 85.67
OWT/${}_{\text{4slices}}$ 91.37 +2.03 91.41 +1.68 87.72 +1.87 87.31 +1.64
Table 3: Performance of semantic segmentation.
Image Segmentation. We observe that reconstructing token groups with OWT preserves clear spatial details and boundary information in Fig. 4. Consequently, a simple thresholding operation on the predictions can generate semantic binary masks, offering direct evidence of each organ-wise token group’s independence and interpretability. Table 3 reports the single-organ segmentation Dice scores (averaged over the liver, kidneys, spleen, and pancreas). Overall, the accuracy of OWT’s organ-specific reconstructions is comparable to that of several dedicated 2D segmentation networks. Moreover, incorporating information from adjacent slices (OWT/${}_{\text{4slices}}$) further improves performance, surpassing the second-best approach by an average margin of 1.805 across four CT and MRI datasets (green number in Table 3). The detailed Dice of each organ are in Appendix 7.
AutoPet Holistic-based Rec. 1 Token Group Rec. 1 Token Group Seg.
L2 $\downarrow$ SSIM-3D $\uparrow$ L2 $\downarrow$ SSIM-3D $\uparrow$ Dice $\uparrow$
MAE/${}_{\text{LPIPS}}$ [27] 3.58e-04 0.9872 - - -
AttU_Net [42] - - - - 87.03
CMUNet [51] - - - - 87.89
OWT 1.12e-04 0.9909 4.07e-04 0.9765 87.54
OWT/${}_{\text{4slices}}$ 9.34e-05 0.9924 3.60e-04 0.9798 89.42
Table 4: Generalization capability of cross-dataset experiments trained on Abdomen1k and tested on AutoPet.
Generalization. Table 4 reports the generalization performance when all methods are trained on Abdomen1k and tested on the AutoPet dataset. Under both reconstruction and segmentation metrics, OWT/${}_{\text{4slices}}$ outperforms dedicated holistic-based approaches. Moreover, it also achieves strong results in single-organ reconstruction, indicating that token groups can be flexibly combined or used individually to mitigate interference from irrelevant organs and background regions. Consequently, the semantic features extracted by OWT generalize more effectively, underscoring its potential as a foundational tokenization framework.
Figure 6: t-SNE visualization of independent token groups on four medical datasets.
4.2.2 Look Deeper into Token Groups
Semantic Information in Token Groups. Fig. 6 illustrates the t-SNE results of how the Organ Collector (with TGR) embeds the background and four organs into the latent space (i.e., five token groups) for each dataset. Within this space, token vectors of the same group type exhibit dense clusters, whereas different token groups remain distinctly separated. This clear inter-group separation highlights the semantic independence of organ-specific token groups in OWT, indicating strong potential for downstream tasks.
Due to the semantic independence of the token group, Fig. 5 shows that our approach not only supports holistic-level retrieval, but also enables organ-level retrieval ($L_{2}$ distance measure). Specifically, we demonstrate top-1, top-2, and maximum-distance retrieval results on both CT and MRI modalities for different token groups. The case indices in the top-left corner of each row indicate how retrieval results vary for different organs of interest. Since experiments in reconstruction and segmentation demonstrate that each token group preserves both the spatial structure and texture details of its corresponding organ, this organ-specific retrieval can be highly beneficial in medical scenarios. For example, it can help radiologists quickly locate similar organ regions across multiple patients, facilitate region-based comparisons of pathology progression, or assist in reference image selection for organ-specific diagnoses.
AutoPet Holistic-based Rec. 1 Token Group Rec.
L2 $\downarrow$ LPIPS $\downarrow$ L2 $\downarrow$ LPIPS $\downarrow$
OWT 1.49e-04 0.0232 3.32e-04 0.0071
OWT w/ Text 1.41e-04 0.0221 3.16e-04 0.0068
Table 5: Reconstruction comparison between OWT w/ and w/o incorporated text prompts.
Figure 7: Loss comparison between OWT w/ and w/o incorporated text prompts.
Multi-Modality Incorporation. In OWT, each organ is embedded into its corresponding SDTG, resulting in highly semantically disentangled and independent representations. Therefore, these representations can be efficiently aligned with or collaborated on with other modality embeddings (e.g., text encoding). Specifically, during OWT training, we use medical prompts and a pretrained CLIP encoder [45] to obtain text encoding of each organ, which are then added element-wise to their matching token groups after the Organ Collector. (More details on prompts and architecture for training OWT w/ text encoding are in Appendix 6.2.2.)
Fig. 7 and Table 5 reveal that simply adding text prompts to organ-wise token groups instantly accelerates loss convergence and provides slight improvements in both holistic- and semantic-based reconstruction. Consequently, OWT shows promise for incorporating the text modality to further enhance organ-specific representations. Such multi-modal integration could unlock new medical applications, such as more robust organ labeling, guided retrieval and generation, and improved interpretability in complex workflows.
4.2.3 Organ-wise Token Groups are more than Organs
CelebAMaskHQ Holistic-based Rec. 1 Token Group Rec.
L2 $\downarrow$ LPIPS $\downarrow$ L2 $\downarrow$ LPIPS $\downarrow$
MAE/${}_{\text{LPIPS}}$ [27] 2.78e-03 0.1157 - -
OWT 2.50e-03 0.1101 2.83e-03 0.0458
Table 6: Performance of OWT on CelebAMaskHQ dataset.
Although OWT and TGR are primarily designed for grouping different semantic components in medical images, they can also be applied to non-medical datasets. As a proof of concept, we demonstrate in Fig. 8 and Table 6 that OWT successfully learns semantically disentangled token groups on the CelebAMaskHQ dataset, capturing and reconstructing each facial region with high spatial and textural fidelity.
Figure 8: Samples of holistic-based and semantic-based human face reconstruction on CelebAMaskHQ.
Figure 9: Semantic-based retrieval of CelebAMaskHQ.
As illustrated in Fig. 9, each token group facilitates semantic-based retrieval for different facial attributes (e.g., hair, skin, lips). In the third column, we conduct an experiment by averaging source token group(s) with the Top1 retrieved token group(s). Notably, this averaged token group(s) can still be reconstructed, exhibiting features that blend and transition between both identities (e.g., overall facial attributes, hair color, skin tone, lip color, mouth openness). These observations imply that token groups extracted by OWT can enable fine-grained editing or hybrid generation of facial attributes without background influence, offering enhanced interpretability, efficiency, and controllability.
5 Conclusion
In this paper, we propose a novel Organ-Wise Tokenization (OWT) framework and Token Group-Based Reconstruction (TGR) paradigm, which disentangle holistic embeddings into semantically disentangled token groups. Our approach not only improves interpretability, generalizability, efficiency and controllability in organ-based analyses, but also shows adaptability to broader scenarios, including natural image domains. By enabling semantic-level token groups, OWT enriches downstream tasks such as semantic-based reconstruction, segmentation and retrieval, offering a new perspective on how representations can be leveraged.
Supplementary Material
Abdomen1k Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
UNetplus [63] 96.85 94.63 91.00 71.05 88.38
AttU_Net [42] 97.09 95.22 93.05 72.00 89.34
TransUnet [9] 96.78 94.65 92.30 71.89 88.91
SwinUnet [5] 96.53 93.68 90.52 65.50 86.56
CMUNet [51] 96.05 94.67 90.84 71.75 88.33
[…]
OWT/${}_{\text{4slices}}$ 96.81 94.84 95.59 78.23 91.37 +2.03
AutoPet Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
UNetplus [63] 96.41 92.56 91.31 76.87 89.29
AttU_Net [42] 96.60 92.86 91.22 78.24 89.73
TransUnet [9] 96.34 93.59 90.65 78.23 89.70
SwinUnet [5] 96.33 92.34 91.11 72.46 88.06
CMUNet [51] 96.56 93.11 90.61 77.65 89.48
[…]
OWT/${}_{\text{4slices}}$ 96.65 93.15 93.96 81.87 91.41 +1.68
Table 7: Performance of semantic segmentation on CT datasets.
Delay Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
UNetplus [63] 93.53 91.85 85.20 65.85 84.11
AttU_Net [42] 94.25 91.08 86.85 65.98 84.54
TransUnet [9] 93.93 90.83 89.07 66.03 84.97
SwinUnet [5] 92.78 90.61 86.35 62.55 83.07
CMUNet [51] 94.02 91.88 87.74 65.64 84.82
[…]
OWT/${}_{\text{4slices}}$ 94.72 92.61 92.23 71.33 87.72 +1.87
PreArtery Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
UNetplus [63] 94.05 92.25 87.26 64.14 84.42
AttU_Net [42] 94.33 92.32 89.07 64.45 85.04
TransUnet [9] 94.33 90.69 86.59 65.56 84.29
SwinUnet [5] 92.91 88.91 86.33 61.48 82.41
CMUNet [51] 94.26 92.58 89.21 66.18 85.56
[…]
OWT/${}_{\text{4slices}}$ 94.74 92.18 92.14 70.17 87.31 +1.64
Table 8: Performance of semantic segmentation on MRI datasets.
Figure 10: Comparison of segmentation performance using direct and indirect approaches.
6 Implementation Details
6.1 Image Preprocessing
For all 3D medical volumes, we standardize the spatial resolution to $2\text{mm}\times 2\text{mm}\times 2\text{mm}$ and apply a center crop of $224\times 224$. For CT datasets, we clip intensity values to the $[-1000,1000]$ Hounsfield Unit (HU) range and normalize them to $[0,1]$. For MRI datasets, we clip intensities to the $[0,99]$ percentile range and normalize them to $[0,1]$. We define five token groups to extract semantic features based on common annotations: background, liver, kidneys, spleen, and pancreas. Each medical image is further standardized in thickness by center-padding or cropping 112 slices based on the anatomical centers of the four organs. For the Abdomen1K dataset [40], the number of cases in the training and test sets is 899:101. For AutoPet [21], the split is 700:200, while for RAOS delay and pre-artery sequences [39], both have a training-to-test ratio of 330:83. All experiments, including OWT, MAE, and segmentation networks, are based on the same preprocessed dataset.
For the 2D face images in CelebAMaskHQ dataset [34], we resize them to $224\times 224$, and define six semantic entities based on annotations (background, face&ear, hair&hat, mouth, nose, and eyes&glasses), as shown in Fig. 8. The CelebAMaskHQ dataset contains 30,000 facial image cases, with 24,000 images allocated for training and 6,000 for testing.
6.2 Architecture Details
6.2.1 Default Architecture of MAE and OWT
In OWT, we adopt the same encoder and decoder architecture as MAE, with the only modification being the replacement of the self-attention blocks [53] with linear self-attention blocks [31]. By default, we use the ViT-base architecture, where the MAE encoder consists of 12 attention blocks, the decoder has 8, and the token embedding dimension is set to 768.
For OWT, we configure the encoder and Token Group Encoder with 6 attention blocks each, while keeping the decoder at 8 blocks. This ensures that the overall model length remains consistent with MAE. Additionally, since each SDTG contains a default of 20 tokens, the computational load of OWT is lower than that of MAE when using fewer than 10 SDTGs. To prevent information leakage during TGR training, we remove the class token from the OWT architecture. For OWT processing 2D images, we use a $16\times 16$ kernel for patch embedding. For 3D input volumes, we apply a $1\times 16\times 16$ 3D kernel for patch embedding.
6.2.2 OWT w/ Text Encoding
For the OWT experiment incorporating text encoding, we design a text prompt inspired by [35]. Specifically, we define the template as "a computerized tomography [category]", where [category] corresponds to "background without the Liver, Kidneys, Spleen, and Pancreas", "of a Liver", "of Kidneys", "of a Spleen", and "of a Pancreas", aligning with the number of token groups. These prompts are fed into a pretrained CLIP text encoder [45] to obtain the corresponding text encoding, resulting in a representation of size $5\times 512$.
To match the token groups, we replicate each text encoding according to the number of tokens per token group (20), leading to a final text encoding dimension of $100\times 512$. Within the OWT framework, a linear transformation layer maps the text encoding from $512$ to $768$, producing a transformed text embedding of size $100\times 768$. This transformed text encoding is then element-wise added to the organ-wise representation $X_{G}$ extracted by the Organ Collector. The combined representation proceeds through the random selection procedure in TGR, while all other components remain consistent with the default OWT settings.
6.3 Experimental Setup
[…]
For OWT with 2D image inputs, we set the batch size to 96. When processing 4-slice medical volumes, the batch size is set to 64. In medical imaging training, each unique input (whether a single slice or a 4-slice volume) is treated as an individual input. For instance, a 3D medical volume with a thickness of 112 slices is considered as 112 separate inputs during training. An epoch is defined as the point when all unique inputs in the dataset have been processed once.
For MAE training, we randomly mask 75% of the tokens. During testing, we adjust the mask ratio to 0.25 for improved performance. When MAE is trained with LPIPS loss, all tokens are preserved for image reconstruction.
All reconstructed images are filtered with a 0.02 pixel-range threshold to reduce background noise. For the interpretability experiments validated by segmentation, we apply a 0.15 threshold to transfer the reconstructed results as semantic masks.
7 Detailed Semantic Segmentation Results
Detailed Results. Tables 7 and 8 provide a more detailed results of those presented in Table 3, showing the Dice scores for each of the four organs across different datasets. From Tables 7 and 8, we observe that compared to dedicated segmentation networks, OWT/${}_{\text{4slices}}$ generally achieves superior performance across all organs. Notably, for smaller organs such as the pancreas, OWT/${}_{\text{4slices}}$ demonstrates a significant improvement, with a Dice score advantage exceeding 5.3 in the best case.
Token Group Connectivity. For semantic segmentation using OWT, we generate masks by processing independently utilized Semantically Disentangled Token Groups (SDTGs) through the Token Group Encoder, AHER, and decoder to reconstruct an image, which is then thresholded to produce a segmentation mask. We refer to this mask as the “direct” mask. Since SDTGs can be flexibly combined, we also explore an alternative approach: subtracting the reconstructed images from all other four SDTGs from the input image and applying a threshold to obtain a predicted mask for the same organ. We define this as the “indirect” mask.
Although using all complementary SDTGs reduces computational efficiency compared to single SDTG usage, it incorporates additional inter- and intra-organ information, potentially enhancing connectivity in the segmentation process. To evaluate whether leveraging more SDTGs improves boundary and shape reconstruction, we conduct comparative experiments. The results, presented in Fig. 10, show that across four datasets, the direct and indirect methods yield largely similar outcomes. This demonstrates that the SDTGs extracted by OWT maintain a high degree of semantic independence, as they accurately capture organ boundaries and shapes even when used independently. These findings align with the t-SNE visualization of SDTG relationships in Fig. 6, further validating their disentangled nature.
Method AutoPet CelebAMaskHQ
L2 $\downarrow$ LPIPS $\downarrow$ L2 $\downarrow$ LPIPS $\downarrow$
MAE/${}_{\text{LPIPS}}$ 2.77e-04 0.0255 2.78e-03 0.1157
OWT 1.49e-04 0.0232 2.50e-03 0.1101
OWT/${}_{\text{VQGAN}}$ 6.89e-05 ↓53.8% 0.0073 ↓68.5% 1.80e-03 ↓43.8% 0.0540 ↓51.0%
Table 9: Performance comparison when altering the Transformer-based encoder and decoder to those of VQGAN.
8 Ablation Studies
For the ablation study, we explore the following key aspects: the flexibility of the encoder-decoder architecture, the number of tokens per SDTG, and the impact of varying the amount of input data in OWT.
Backbone Alteration. We investigate the effect of replacing the default ViT-based encoder-decoder architecture in OWT (adopted from MAE) with the encoder and decoder from VQGAN. Specifically, we configure the VQGAN bottleneck dimensions to $14\times 14$ for a $224\times 224$ input image, ensuring that the number of latent tokens (196) remains consistent with the ViT-based setting. As shown in Table 9, despite the increased computational cost of VQGAN’s hybrid CNN and self-attention structure, it achieves significant performance gains over the transformer-based encoder-decoder. On two (medical and facial) datasets, it improves both L2 and LPIPS evaluation metrics by 43.8%–68.5%, demonstrating the potential for further enhancements in OWT and highlighting its architectural flexibility.
Figure 11: Ablation study on the number of tokens in each token group when trained on AutoPet.
Figure 12: Ablation study on OWT performance when utilizing different percentages of the dataset (AutoPet).
Number of Tokens per SDTG. The number of tokens within each SDTG determines its capacity to store organ-specific information. To evaluate this, we experiment with SDTG token counts of 5, 10, 20, and 40, assessing their impact on both generation and segmentation tasks. For the generation task, we measure full-image reconstruction using all SDTGs and single-organ reconstruction using only one SDTG, evaluated by L2, LPIPS, and SSIM-3D metrics. For segmentation, we assess the Dice score of single-organ masks.
As shown in Fig. 11, increasing the token count from 10 to 20 provides a slight performance improvement, while further increasing it from 20 to 40 leads to a more significant boost. This suggests that OWT has considerable scalability and potential as a foundational semantic-level tokenization framework, capable of further enhancements with increased token capacity.
Data Percentage. Fig. 12 illustrates the performance of the OWT framework on the AutoPet dataset when trained with 20%, 60%, and 100% of the data. We observe a significant improvement in performance when increasing the dataset size from 20% to 60%, whereas the gain from 60% to 100% is relatively marginal. This suggests that OWT efficiently learns meaningful representations with a moderate amount of data, indicating its strong data efficiency and the potential for effective generalization even with limited training samples.

### v2

@@
 Abstract
-Recent advances in representation learning often rely on holistic, black-box embeddings that entangle multiple semantic components, limiting interpretability and generalization.
-These issues are especially critical in medical imaging.
+Recent advances in representation learning often rely on holistic embeddings that entangle multiple semantic components, limiting interpretability and generalization.
+These issues are especially critical in medical imaging, where downstream tasks depend on anatomically interpretable features.
 To address these limitations, we propose an Organ-Wise Tokenization (OWT) framework with a Token Group-based Reconstruction (TGR) training paradigm.
-Unlike conventional approaches that produce holistic features, OWT explicitly disentangles an image into separable token groups, each corresponding to a distinct organ or semantic entity.
-Our design ensures each token group encapsulates organ-specific information, boosting interpretability, generalization, and efficiency while allowing fine-grained control in downstream tasks.
-Experiments on CT and MRI datasets demonstrate the effectiveness of OWT in not only achieving strong image reconstruction and segmentation performance, but also enabling novel semantic-level generation and retrieval applications that are out of reach for standard holistic embedding methods.
-These findings underscore the potential of OWT as a foundational framework for semantically disentangled representation learning, offering broad scalability and applicability to real-world medical imaging scenarios and beyond.
-[^1] Code is available at https://github.com/SifanSong/OWT
+Unlike conventional approaches, OWT explicitly disentangles an image into separable token groups, each corresponding to a distinct organ or semantic entity.
+Our design ensures each token group encapsulates organ-specific information, boosting interpretability, generalization, and efficiency while enabling fine-grained control for targeted clinical applications.
+Experiments on CT and MRI datasets demonstrate OWT’s power: it not only achieves strong performance on standard tasks like image reconstruction and segmentation, but also unlocks novel, high-impact clinical capabilities including organ-specific tumor identification, organ-level retrieval and semantic-level generation, without requiring any additional training.
+These findings underscore the potential of OWT as a foundational framework for semantically disentangled representation learning, offering broad scalability and a new perspective on how representations can be leveraged.
 1 Introduction
-Recent years have witnessed significant advances in representation learning of visual data for both natural and medical images.
-The typical representation learning, driven by self-supervised and task-guided training frameworks, is classified as holistic embedding-based representation learning since it learns embeddings that capture all information for an entire image [54, 22, 58] (Fig.1-a).
-Similar trends appear in the medical imaging domain, where holistic representations have also been extensively explored for tasks like disease diagnosis[49], organ segmentation [46, 5, 25], and anatomical structure analysis [41, 59].
-Although the holistic embeddings leveraged by these studies have been proven effective, they remain globally entangled and mixed in a black-box latent space, combining information about multiple organs and background.
-This fusion decreases efficiency and poses significant challenges for organ-focused analyses, disease diagnosis, or radiological interpretation, where fine-grained interpretability and a semantically targeted focus are critically important [48, 6, 58].
 Figure 1: Objective comparison of (a) holistic embedding-based representation learning, (b) disentangled representation learning, and (c) our OWT framework.
@@
 These token groups can be leveraged separately or in combination for downstream analyses, such as semantic-level reconstruction, segmentation, and retrieval.
-In parallel, disentangled representation learning (DRL) has been studied to decompose data into semantically meaningful factors of variation, with each factor ideally corresponding to an independent aspect of the underlying data [58] (Fig.1-b).
-By isolating latent factors (e.g., shape, color, orientation) in natural images [4, 12, 61, 64] or modality attributes in medical images [44, 8], DRL methods aim to improve explainability and enable controllable feature manipulation.
-However, most conventional DRL approaches rely on purely data-driven or perturbation-driven strategies, requiring extensive modeling assumptions or specialized regularizers to ensure each factor corresponds to a semantically coherent concept [38, 58].
-Consequently, existing DRL approaches face notable challenges, including limited alignment with real-world semantics, weak spatial correspondence, and poor scalability efficiency.
-Despite these successes, existing holistic embedding–based strategies and DRL approaches still have notable gaps, particularly in medical imaging, where high-level interpretability and organ-specific analysis are crucial (e.g., anatomical analysis and surgical planning).
-Specifically, existing methods suffer from (a) a lack of interpretability and generalizability, as multiple semantic cues are often merged without alignment to real-world concepts, making it difficult to isolate relevant structures from confounding factors (e.g., background variations); and (b) limited efficiency and controllability, since these approaches process entire or masked representations rather than focusing only on organ-level features, restricting the ability to selectively manipulate or analyze distinct semantic factors.
-These limitations motivate us to revisit the design of representation learning, aiming for explicit organ- (or part-) level tokenization to tackle them.
-In this paper, we propose a foundational Organ-Wise Tokenization (OWT) framework and a novel Token Group-based Reconstruction (TGR) training paradigm.
-Specifically, in addition to a standard encoder-decoder architecture, such as Masked Autoencoders (MAE) [27], our OWT framework incorporates two key novel modules, Organ Collector and Adaptive Holistic Embedding Restorer (AHER), inspired by prior structures for extracting semantic features [47, 18].
-The Organ Collector integrates holistic embeddings into semantically meaningful token groups, while AHER reconstructs holistic representations from any number of token groups.
-Together, these modules enable a structured transition from holistic $\rightarrow$ semantic $\rightarrow$ holistic embeddings, facilitating both organ-specific processing and reconstruction.
-Meanwhile, our proposed TGR training paradigm ensures that the extracted semantic embeddings are fully functional as Semantically Disentangled Token Groups (SDTGs) by explicitly correlating each token group to its corresponding organ.
-In this case, each SDTG encapsulates only organ-specific information and is capable of being utilized independently or in any combination with other SDTGs as practically needed (Fig.
+Representation learning has advanced rapidly in recent years, with growing impact across both natural and medical imaging domains, enabling more effective feature extraction and improved performance across a wide range of downstream tasks.
+Building on this progress, a variety of self-supervised and task-guided training frameworks have been widely adopted, typically relying on holistic embeddings that capture aggregated information for an entire image without explicitly separating semantically distinct components.
+[voulodimos2018deep, geirhos2020shortcut, wang2024disentangled] (Fig.1-a).
+Similar trends appear in the medical imaging domain, where holistic representations have also been extensively explored for tasks like disease diagnosis [shen2017deep], organ segmentation [ronneberger2015u, cao2022swin, hatamizadeh2021swin], and anatomical structure analysis [mortada2023segmentation, weston2019automated].
+Although holistic embeddings have shown strong performance, they remain globally mixed within a latent space, entangling information across multiple organs and background regions without explicit semantic separation.
+This mixed representation poses significant challenges for where fine-grained interpretability and a semantically targeted focus are critically important [schutte2021using, cetin2023attri, wang2024disentangled].
+On the other hand, disentangled representation learning (DRL) aims to decompose data into semantically meaningful factors, each capturing an independent aspect of the underlying variation [wang2024disentangled] (Fig.1-b).
+DRL isolates latent factors (e.g., shape, color, orientation) in natural images [burgess2018understanding, chen2016infogan, xiao2017dna, zhu2018visual] or modality attributes in medical images [ouyang2021representation, chartsias2019disentangled], thereby improving explainability and enabling controllable feature manipulation.
+However, most existing DRL approaches rely on purely data-driven or perturbation-driven strategies, requiring extensive modeling assumptions or specialized regularizers to ensure each factor corresponds to a semantically coherent concept [locatello2019challenging, wang2024disentangled].
+Consequently, existing DRL approaches suffer from limited alignment with real-world semantics, weak spatial correspondence, and poor scalability (More details on Related Work are provided in Appendix A).
+While DRL offers a promising direction, both holistic embeddings and existing disentanglement approaches remain limited in medical imaging scenarios, where organ-level interpretability and analysis are important (e.g., anatomical analysis and surgical planning).
+These limitations manifest in two key aspects: (a) a lack of interpretability and generalizability, as multiple semantic cues are often entangled without alignment to real-world concepts, making it difficult to isolate relevant structures from confounding factors (e.g., background); and (b) limited efficiency and controllability, as these approaches operate on fused representations rather than isolated organ-level features, restricting the ability to selectively manipulate or analyze distinct semantic factors.
+These limitations motivate the development of our novel representation learning framework that explicitly tokenizes images at the organ or semantic level.
+In this paper, we propose a foundational Organ-Wise Tokenization (OWT) framework and a novel Token Group-based Reconstruction (TGR) training paradigm (Fig.
 1-c).
-Therefore, our design offers key advantages by addressing the limitations of existing approaches: (i) improved interpretability and generalizability, as each token group is explicitly associated with a specific organ or semantic content, enabling the isolation of target structures while minimizing confounding factors; and (ii) enhanced efficiency and controllability, as OWT processes only the relevant token groups for a given task, reducing computational load and allowing for semantic-based, fine-grained manipulation and adaptation without introducing unrelated features.
+The OWT framework extends a standard encoder-decoder architecture, such as Masked Autoencoders (MAE) or VQGAN, by incorporating two specialized modules: Organ Collector and the Adaptive Holistic Embedding Restorer (AHER), which are inspired by prior work on semantic tokenization [ryoo2021tokenlearner, esser2021taming, fayyaz2022adaptive].
+The Organ Collector integrates holistic embeddings into semantically meaningful Token Group (TG), while AHER enables to reconstruct holistic representations from any number of token groups.
+This design enables a transition pathway from holistic $\rightarrow$ semantic $\rightarrow$ holistic embeddings, facilitating both organ-specific processing and reconstruction.
+To ensure that the extracted semantic embeddings function as true Semantically Disentangled Token Groups (SDTGs), the proposed TGR training paradigm explicitly enforces correspondence between each token group and its target organ.
+These components constitute a unified framework that encodes organ-specific semantics into modular and reusable TGs, enabling task-adaptive control and restricting computation to relevant semantic regions without additional training, thereby enhancing interpretability and eliminating redundant processing.
 Figure 2: Overall architecture of Organ-Wise Tokenization (OWT).
@@
 Figure 3: Comparison of the overall process of holistic embedding-based and token group-based reconstruction.
-We evaluate the performance of the organ-wise embeddings (i.e., SDTGs) extracted by OWT across multiple downstream tasks (such as image reconstruction, segmentation and retrieval).
-Across multiple modalities (CT, two MRI sequences, and a facial dataset), our method achieves better and more robust results compared to baselines (i.e., MAE [27] and dedicated semantic segmentation networks [63, 42, 9, 5, 51]) and is able to accomplish tasks that these traditional methods cannot, such as semantic-level generation, retrieval and embedding combination.
-This proves the concept of our foundational semantically disentangled tokenization framework, enabling us to bring more potential solutions for real-world applications.
+We evaluate the organ-wise embeddings on downstream tasks (e.g., image reconstruction and segmentation).
+Across multiple modalities (CT, two MRI sequences, and a facial dataset), our method consistently outperforms strong baselines, including VAE [kingma2013auto], MAE [he2022masked], and state-of-the-art task-specific segmentation architectures such as UNet++ [zhou2018unet++], AttU_Net [oktay2018attention], TransUNet [chen2021transunet], Swin-UNet [cao2022swin], and CMUNet [tang2023cmu].
+Beyond these general tasks, OWT unlocks capabilities that conventional models cannot provide: organ-level tumor identification and retrieval, semantic-level generation, and flexible combination of organ embeddings to adapt to heterogeneous downstream requirements.
+This demonstrates that the semantic disentanglement enabled by our OWT framework can be leveraged for broader real-world applications, offering a new perspective on how representations can be constructed and utilized.
 In summary, our contributions are listed as follows:
-  • We propose a foundational framework and a novel training paradigm, OWT and TGR, which encourage the network to encapsulate organ-level features into semantically disentangled token groups (SDTGs) rather than entangled embeddings with a black-box nature.
-  • The proposed SDTGs are highly condensed with rich information on each target organ, which increases the interpretability and generalization of latent embeddings.
-  • The token groups can be used separately or in combination and are processed with improved efficiency and controllability for downstream tasks and further analysis.
-  • Experiments demonstrate that our method offers significant scalability.
-It can adaptively integrate semantic information based on various requirements and incorporate other modalities (e.g., text encoding), highlighting its potential for broader applications beyond medical imaging.
-2 Related Work
-2.1 Holistic Embedding-based Representation Learning
-Holistic embedding-based approaches map entire 2D/3D inputs into a single latent representation, aiming to capture global image features [54, 22, 58].
-These methods typically follow either self-supervised or task-guided strategies.
-In self-supervised learning, contrastive techniques like SimCLR [10], SimSiam [11], and BYOL [24] enforce consistency between augmented views, while predictive approaches such as CPC [43] and RotNet [23] reconstruct missing or permuted portions.
-Recently, masked image modeling (e.g., MAE [27] and BEiT [3]) has emerged as an effective approach by requiring models to infer masked regions.
-Meanwhile, task-guided frameworks focus on extracting features that directly benefit a target task, typically resulting in holistic embeddings.
-Many classic neural networks, such as ResNet [26], DenseNet [29], and Vision Transformer [16], and their variants incorporate classification, regression, or segmentation objectives.
-In medical imaging, holistic representations have been widely used for self-supervised feature learning in CT [13, 32], MRI [7, 33], and X-ray [52, 60].
-Additionally, classical supervised pipelines (e.g., U-Net [46] and its variants [63, 9, 5, 25]) rely on a global bottleneck to embed entire images or volumes, demonstrating strong performance in organ segmentation and disease detection.
-Although such approaches achieve notable success, they often encode multiple semantic factors, such as multiple organs and the surrounding background, into an entangled representation with limited interpretability and fine-grained control [48, 6, 58].
-2.2 Disentangled Representation Learning
-In contrast to holistic embedding-based approaches, various disentangled representation learning (DRL) studies have pursued a dimension-wise strategy for two main objectives.
-First, to derive more robust holistic embeddings with improved generalizability [14, 1, 15, 56, 20, 50, 58], and second, to explicitly decompose images into latent factors, each corresponding to distinct and semantically meaningful attributes, in order to improve explainability [4, 12, 61, 64].
-In medical imaging, the motivations for DRL are especially strong: improved interpretability, controllability, and robustness to confounding factors are crucial in clinical settings [37, 55].
-When transitioning DRL to medical imaging, researchers seek to decouple anatomical properties (e.g., organ size or shape) [57, 30], modality characteristics (e.g., CT and MRI sequences) [44, 8], and pathological factors (e.g., tumor presence or severity) [36, 28], aiming to isolate distinct clinical features.
-Despite notable progress, data- and perturbation-driven DRL approaches can be sensitive to training objectives, architecture choices, and assumptions about independent factors [38, 56, 58].
-In existing studies, disentangled dimensions may not align precisely with real anatomical or pathological concepts, and spatial correspondence often remains weak when images are globally encoded.
-Consequently, there has been a growing emphasis on methods that combine disentanglement principles with controlled grouping or tokenization schemes to encapsulate semantically meaningful subsets of features.
+  • We introduce OWT, the first organ-wise foundational framework that, paired with a novel TGR training paradigm, encapsulates organ-specific semantics into semantically disentangled token groups (SDTGs) within a single training process.
+  • The learned SDTGs are highly condensed yet rich in information for each target organ, enhancing the interpretability and generalizability of the latent embeddings.
+  • The SDTGs can be selectively combined and reused, yielding improved efficiency and fine-grained controllability.
+This unlocks unique organ-level downstream tasks, such as organ-specific retrieval and identification, which are highly demanded in the medical imaging domain.
+  • Our experiments demonstrate significant scalability, as OWT adaptively disentangles semantic information across diverse modalities and domains.
+This includes a real-world facial dataset, highlighting its potential for applications beyond medical imaging.
 2 Method
-In this paper, we present a novel OWT framework and a Token Group-based Reconstruction (TGR) training paradigm, which embeds each organ into its corresponding semantically disentangled token group (SDTG).
-In this case, these organ-specific embeddings can be utilized independently or in combination for downstream tasks.
-3.1 Overall OWT Framework
-As illustrated in Fig.
-2, our proposed OWT framework comprises five components: an encoder, Organ Collector, Token Group Encoder, Adaptive Holistic Embedding Restorer (AHER), and a decoder.
+We propose Organ-Wise Tokenization (OWT) with Token Group-based Reconstruction (TGR): OWT encodes each organ into a semantically disentangled token group (SDTG) that can be used alone or composed for downstream tasks.
+Overall, the proposed OWT framework comprises five components (Fig.
+2): an encoder, Organ Collector, Token Group Encoder, Adaptive Holistic Embedding Restorer (AHER), and a decoder.
 The encoder maps input images to holistic embeddings, while the Organ Collector and TGR enable the target-organ features to be extracted and stored in specific embeddings (i.e., SDTGs).
@@
 Finally, the decoder reconstructs an image of the same size as the input.
-The encoder and decoder in OWT follow standard holistic representation learning architectures, so they can be Transformer-based or (fused) CNN-based, as in 2D/3D MAE [27, 19] or 2D/3D VQGAN [17, 19].
+2.1 Encoder and Decoder in OWT
+The encoder and decoder in OWT follow standard holistic representation learning architectures, so they can be Transformer-based or (fused) CNN-based, as in 2D/3D MAE [he2022masked, feichtenhofer2022masked] or 2D/3D VQGAN [esser2021taming, feichtenhofer2022masked].
 This design choice allows the framework to be applied to either 2D or 3D images with minimal modification.
@@
 2.2 Organ Collector
-To effectively extract semantically meaningful features from the holistic embedding $X_{H}$, we introduce a novel Organ Collector module.
-A key component of this process is the token-wise attention matrix $A_{t}$, which dynamically assigns importance to different spatial locations or tokens, ensuring that each token group captures the most relevant semantic features.
-To achieve this, let $X_{H}\in\mathbb{R}^{(dhw)\times c_{e}}$ be reshaped to $\mathbb{R}^{c_{e}\times d\times h\times w}$ before being fed into the Organ Collector (Fig.
+To effectively extract semantically meaningful features from holistic embedding $X_{H}$, the Organ Collector module dynamically assigns importance to different spatial locations or tokens using a token-wise attention matrix $A_{t}$.
+Let $X_{H}\in\mathbb{R}^{(dhw)\times c_{e}}$ be reshaped to $\mathbb{R}^{c_{e}\times d\times h\times w}$ before being fed into the Organ Collector (Fig.
 2).
@@
 Concretely, $\alpha$ projects the $(c_{e}\times d)$-dimensional input to $N_{g}$, then applies Flatten and Softmax to derive a token-wise attention matrix.
-This attention matrix is multiplied by $\gamma(X_{H})$ (also flattened) to obtain $X_{G}\in\mathbb{R}^{N_{g}\times c_{e}}$.
-Abdomen1k Holistic-based Rec. 1 Token Group Rec.
-L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$ L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$
-MAE [27] 6.74e-04 0.0684 0.9688 - - -
-MAE/${}_{\text{LPIPS}}$ 6.19e-04 0.0403 0.9722 - - -
-OWT 4.16e-04 0.0390 0.9734 3.31e-04 0.0073 0.9795
-OWT/${}_{\text{4slices}}$ 3.69e-04 0.0371 0.9755 2.68e-04 0.0069 0.9838
-AutoPet Holistic-based Rec. 1 Token Group Rec.
-L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$ L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$
-MAE [27] 2.86e-04 0.0452 0.9799 - - -
-MAE/${}_{\text{LPIPS}}$ 2.77e-04 0.0255 0.9843 - - -
-OWT 1.49e-04 0.0232 0.9881 3.32e-04 0.0071 0.9778
-OWT/${}_{\text{4slices}}$ 1.13e-04 0.0179 0.9908 2.68e-04 0.0063 0.9827
-Table 1: Performance of image reconstruction on CT datasets.
-Abdomen1k Holistic-based Rec. 1 Token Group Rec.
-L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$ L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$
-MAE [27] 1.37e-03 0.0465 0.9686 - - -
-MAE/${}_{\text{LPIPS}}$ 1.36e-03 0.0389 0.9690 - - -
-OWT 7.65e-04 0.0256 0.9851 5.96e-04 0.0088 0.9762
-OWT/${}_{\text{4slices}}$ 5.81e-04 0.0214 0.9885 5.04e-04 0.0081 0.9801
-PreArtery Holistic-based Rec. 1 Token Group Rec.
-L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$ L2 $\downarrow$ LPIPS $\downarrow$ SSIM-3D $\uparrow$
-MAE [27] 1.67e-03 0.0438 0.9701 - - -
-MAE/${}_{\text{LPIPS}}$ 1.74e-03 0.0375 0.9672 - - -
-OWT 8.66e-04 0.0248 0.9851 5.29e-04 0.0089 0.9759
-OWT/${}_{\text{4slices}}$ 7.07e-04 0.0218 0.9876 4.47e-04 0.0082 0.9801
-Table 2: Performance of image reconstruction on MRI datasets.
+This attention matrix is multiplied by $\gamma(X_{H})$ (also flattened and transposed) to obtain $X_{G}\in\mathbb{R}^{N_{g}\times c_{e}}$.
+This ensures that each token group captures the most relevant semantic features.
+Figure 4: Samples of holistic-based and semantic-based medical image reconstruction from an MRI dataset.
 2.3 Token Group-based Reconstruction (TGR)
-The TGR training paradigm provides direct supervision to token groups, ensuring their semantic disentanglement.
+The TGR training paradigm provides direct supervision of semantic disentanglement to token groups.
 By enforcing accurate organ-wise feature extraction and reconstruction through the token group-level loss $\mathcal{L}_{T}$ (Fig.
@@
 Consequently, the extracted semantic tokens serve as semantically disentangled token groups (SDTGs).
-Through TGR, the Organ Collector adaptively extracts organ-specific information into its corresponding SDTG, assigning each group to a particular organ.
+Through TGR, the Organ Collector adaptively extracts organ-specific information into its corresponding SDTG, assigning each token group to a particular organ.
 Definition of Token Group.
-Given a dataset $D$, we have its ground-truth annotations for $g$ organs.
+Given a dataset $D$, we have its ground-truth annotations for $g$ semantic group (such as organs in medical image).
 The remaining regions are treated as background and assigned to a single additional token group.
@@
 Random Selection Procedure.
-To manage these token groups so that they are responsible for each organ during training, we adopt a two-stage random selection procedure.
+To establish and preserve the correspondence between token groups and their target organs during training, we apply a two-stage random selection.
 First, randomly determine the number of retained groups: Let $\Omega\sim\mathcal{U}(0,1)$ be a uniform random variable on $[0,1]$.
@@
 As illustrated in Fig.
-3, TGR departs from conventional holistic embedding-based reconstruction methods that rely on the latent space for the entire image [2, 27, 19, 17] (Fig.
+3, TGR departs from conventional holistic embedding-based reconstruction methods that rely on the latent space for the entire image [bank2023autoencoders, he2022masked, feichtenhofer2022masked, esser2021taming] (Fig.
 3-a).
 By selectively preserving certain token groups, our approach avoids fusing diverse organ features and does not require masked tokens, resulting in improved interpretability, efficiency, and controllability during both reconstruction and subsequent tasks.
-After that, we employ a Token Group Encoder ($\mathbf{Enc}_{G}$) composed of six multi-head linear self-attention blocks [31].
+After that, we employ a Token Group Encoder ($\mathbf{Enc}_{G}$) composed of six multi-head linear self-attention blocks [katharopoulos2020transformers].
 Let $\tilde{X}^{\prime}_{G}=\mathbf{Enc}_{G}(\tilde{X}_{G})\in\mathbb{R}^{\tilde{N}_{g}\times c_{e}}$ represent the further contextual modeling that captures inter- and intra-organ relationships of the retained token groups.
@@
 Although $\tilde{N}_{g}$ varies per input, this operation preserves organ-specific features while ensuring $X^{\prime}_{H}$ remains a fixed $(dhw)\times c_{e}$ shape, thereby maintaining consistency with the decoder.
-Finally, we pass $X^{\prime}_{H}$ through a standard decoder [27] ($\mathbf{Dec}$) and unpatchify the output, reconstructing the image $\hat{I}_{T}=\mathbf{Dec}(X^{\prime}_{H})\in\mathbb{R}^{D\times H\times W\times C}$.
+Finally, we pass $X^{\prime}_{H}$ through a standard decoder [he2022masked] ($\mathbf{Dec}$) and unpatchify the output, reconstructing the image $\hat{I}_{T}=\mathbf{Dec}(X^{\prime}_{H})\in\mathbb{R}^{D\times H\times W\times C}$.
 As mentioned in Sec.
-3.3, our objective is to reduce the discrepancy between the target image $I_{target}$ and the reconstructed image $\hat{I}_{T}$.
-We combine an L2 term with the LPIPS perceptual loss [62]
+2.3, our objective is to reduce the discrepancy between the target image $I_{target}$ and the reconstructed image $\hat{I}_{T}$.
+We combine an L2 term with the LPIPS perceptual loss [zhang2018unreasonable]
 $\mathcal{L}_{T}=\|I_{target}-\hat{I}_{T}\|_{2}^{2}\ +\mathcal{L}_{LPIPS}(I_{target},\hat{I}_{T})\;$ (5)
 to preserves low-level fidelity while capturing high-level perceptual similarity, thereby guiding the network to generate more realistic texture and coherent organ regions.
-4 Experiments
-Figure 4: Samples of holistic-based and semantic-based medical image reconstruction from an MRI dataset.
-4.1 Datasets and Experimental Setup
-Datasets.
-We conduct experiments on four 3D medical datasets: two CT sets (Abdomen1k [40] and the CT volumes from AutoPet [21]) and two synthetic MRI sequences (delay and pre-artery) from RAOS dataset [39].
+Tasks Organ-specific Organ-level Pre-defined Any Comb. Segment-
+Identification Retrieval Rec. Rec. ation
+Rec. Nets
+Seg. Nets
+[…]
+Table 1: Downstream task capabilities after one-time training.
+3 Datasets and Experimental Setup
+We conduct experiments on four 3D medical datasets: two CT sets (Abdomen1k [ma2021abdomenct] and the CT volumes from AutoPet [Gatidis2022]) and two synthetic MRI sequences (delay and pre-artery) from RAOS dataset [luo2024rethinking].
 These multimodal and multi-sequence datasets validate the effectiveness and robustness of the OWT framework in medical imaging.
-To further validate that OWT can generalize to different semantic groupings beyond medical scenarios, we employ the CelebAMaskHQ dataset [34].
-More details on the preprocessing steps are in Appendix 6.1.
-Experimental Setup.
+To further validate that OWT can generalize to different semantic units beyond medical scenarios, we employ the CelebAMaskHQ dataset [lee2020maskgan].
+More details are in Appendix B.1, including image preprocessing and data splitting.
 All experiments are conducted on six Nvidia H100 GPUs.
-The hyperparameters, optimizer and scheduler settings follow MAE [27].
-We compare OWT against a baseline MAE due to the similar encoder-decoder structure.
-In OWT, the default token number for each token group is 20.
-For processing 3D volumes (4 slices by default), we adopt a 3D version of MAE [19].
-For fairness, we add LPIPS loss [62] to MAE to improve reconstruction quality.
-We also explore replacing the encoder-decoder pair with that of VQGAN [17].
-Further details of architectures and hyperparameters are in Appendices 6.2 and 6.3, and ablation studies of default settings are in Appendix 8.
-4.2 Results
-Our OWT framework not only yields highly condensed representations containing rich information but also provides semantically meaningful and disentangled token groups.
-In this section, we design extensive experiments to evaluate OWT’s performance on several downstream tasks, including reconstruction, segmentation, generalization, and retrieval (Sec.
-4.2.1).
-Notably, our findings indicate that, in contrast to conventional holistic-based methods, OWT can focus on semantic-level tasks and analyses, delivering superior interpretability and controllability.
-Next, we investigate the characteristics of these semantically disentangled token groups and explore their potential to incorporate other modalities (Sec.
-4.2.2).
-We then demonstrate that the OWT framework can be employed in real-world applications (Sec.
-4.2.3).
-Additionally, we provide ablation studies and explore more architectures of the OWT framework in Appendix 8.
-4.2.1 Downstream Analysis
-Image Reconstruction.
-We evaluate the reconstruction capability of OWT across four datasets (two CT and two MRI) [40, 21, 39].
-Tables 1 and 2 show that OWT outperforms MAE on L2, LPIPS, and 3D SSIM metrics in holistic-based reconstruction tasks.
-Moreover, leveraging a patch-embedding module that tokenizes four slices for 3D volumes generates better results than the 2D OWT variant.
-We further explore a semantic-based reconstruction scenario by retaining only a single organ-wise token group at a time, feeding it through AHER and the decoder to generate an organ-specific image.
-As reported in Tables 1 and 2, using a larger number of slices in CT and MRI volumes leads to more robust disentangled token groups, thereby improving semantic-based reconstruction quality.
+For fairness, we also add LPIPS loss [zhang2018unreasonable] to VAE and MAE to improve the quality of reconstruction.
+Further details of the architecture descriptions and hyperparameters are in Appendix B.2 and B.3.
+Methods Enc. OC. TG Enc. AHER Dec. Total
+MAE 16.66 - - - 11.11 27.77
+[…]
+Table 2: Comparison of computation costs (GFLOPs).
+Methods Disentangled Used Tumor Identification
+Organ Info. Token Num. Liver Kidneys
+MAE 196 73.5 74.7
+[…]
+Table 3: Performance for the tumor identification tasks.
+4 Results
+In this section, we design extensive experiments to explore OWT’s unique clinical impacts (Sec.
+4.2), and to evaluate performance on several downstream tasks (Sec.
+4.3).
+Next, we dive into SDTGs and investigate their semantic characteristics (Sec.
+7).
+We also explore the potential of our OWT to address the common issue of limited medical annotations (Sec.
+4.5).
+Finally, we demonstrate that the OWT framework can also be employed in natural image applications (Sec.4.6), and provide ablation studies in Appendix D to discuss several important hyperparameters of OWT, including encoder/decoder alterations, different token number per organ, and data percentage used for training.
+Abdomen1k Holistic Rec. 1 Token Group Rec.
+[…]
+[…]
+MAE 6.74e-04 0.0684 0.9688
+[…]
+[…]
+[…]
+AutoPet Holistic Rec. 1 Token Group Rec.
+[…]
+[…]
+MAE 2.86e-04 0.0452 0.9799
+[…]
+[…]
+[…]
+Table 4: Performance of image reconstruction on CT datasets.
+Delay Holistic Rec. 1 Token Group Rec.
+[…]
+[…]
+MAE 1.37e-03 0.0465 0.9686
+[…]
+[…]
+[…]
+PreArtery Holistic Rec. 1 Token Group Rec.
+[…]
+[…]
+MAE 1.67e-03 0.0438 0.9701
+[…]
+[…]
+[…]
+Table 5: Performance of image reconstruction on MRI datasets.
+4.1 Advances in Design and Architecture
+Due to our proposed TGR training paradigm, the OWT framework produces semantically disentangled token groups (SDTGs) rich in organ-specific information.
+Leveraging the AHER module, these SDTGs can be individually or flexibly combined and restored from semantic embeddings into their original spatial dimensions.
+Consequently, as shown in Table 1, OWT requires only one-time training to enable multiple downstream tasks.
+This cross-task efficiency and controllability significantly distinguish OWT from conventional networks with entangled embeddings, which cannot achieve such flexibility.
+OWT is also computationally efficient.
+As shown in Table 2, OWT requires only 24.01, notably fewer than MAE’s 27.77 GFLOPs.
+This efficiency stems from (1) OWT’s semantic tokens containing more highly condensed information than the traditional entangled holistic tokens, and (2) the lightweight design of the Organ Collector (0.25 GFLOPs) and AHER (0.074 GFLOPs) modules.
+These designs also offer substantial potential.
+Specifically, we explore adaptively distributing different token numbers to SDTGs (e.g., based on organ volumes) in Appendix D.3.
+Since Organ Collector and AHER are encoder-decoder independent, we further investigate notable performance gains by replacing Transformer-based structures with VQGAN-based alternatives in Appendix D.1.
+Figure 5: Semantic-based organ-level retrieval of MRI images.
+4.2 Unique Clinical Impacts
+Organ-Specific Tumor Identification.
+In contrast to conventional models with entangled holistic embeddings, OWT extracts organ-specific information into distinct, semantically disentangled token groups (SDTGs).
+This structure enables the efficient use of only relevant tokens for highly specific downstream tasks.
+We demonstrate this on a classification task derived from tumor labels in the Abdomen1k dataset.
+As shown in Table 3, OWT leverages only 20 organ-specific tokens (i.e., for liver or kidneys) yet achieves significantly better performance than the holistic embedding-based MAE, which utilizes 196 tokens.
+This results in accuracy improvements of 6.2% (79.7 vs.
+73.5) and 7.1% (81.8 vs.
+74.7) for identifying tumors in the liver and kidneys, respectively, highlighting OWT’s ability to create condensed yet highly informative representations for targeted clinical tasks.
+Organ-Level Retrieval.
+Due to their semantic independence, our token groups enable organ-level retrieval ($L_{2}$ distance measure), a capability holistic embedding-based models cannot perform (Fig.
+5).
+We demonstrate top-1, top-2, and maximum-distance retrieval results on the MRI modality (CT examples are in Appendix F), where case indices in the top-left corner show results for different organs of interest.
+Since each token group preserves its organ’s spatial structure and texture details, this targeted retrieval is highly beneficial in medical scenarios.
+For example, it can help radiologists quickly locate similar organ regions across patients without being influenced by other organs, facilitate region-based pathology comparisons, or assist in selecting reference images for organ-specific diagnoses.
+4.3 Downstream Analysis
+[…]
+We further explore semantic-based reconstruction by retaining a single organ-wise token group, which is fed through AHER and the decoder to generate an organ-specific image.
+As reported in Tables 4 and 5, using more slices leads to more robust token groups and improves semantic reconstruction quality.
 Fig.
-4 illustrates MRI cases (rich in textural detail).
-Even with LPIPS added, MAE suffers from noticeable detail loss, whereas OWT (all token groups retained) produces sharper and clearer details.
-Additionally, we can freely combine the token groups to reconstruct only the content of interest.
-In the zoom-in rows of Fig.
-4, the liver/kidney reconstruction appears nearly identical whether we use only the liver/kidney token group or combine it with others, suggesting that OWT’s organ token groups are highly semantically independent.
-This property offers better interpretability as well as controllability, enabling reconstruction without unnecessary entanglement of unrelated organ features.
-Figure 5: Semantic-based retrieval of CT/MRI images.
-Methods CT MRI
+4 illustrates this on MRI cases: even with LPIPS added, MAE suffers from noticeable detail loss (red arrows), whereas OWT (all token groups retained) produces sharper, clearer details.
+Additionally, token groups can be freely combined to reconstruct only content of interest.
+The zoom-in rows of Fig.
+4 show that the liver/kidney reconstruction appears nearly identical whether using only its specific token group or combining it with others, suggesting OWT’s organ token groups are highly semantically independent.
+This property enhances interpretability and controllability, enabling reconstruction without unnecessary entanglement of unrelated organ features.
+1- / 4-slice CT MRI
 Abdmen1k AutoPet Delay PreArtery
-UNetplus [63] 88.38 89.29 84.11 84.42
-AttU_Net [42] 89.34 89.73 84.54 85.04
-TransUnet [9] 88.91 89.70 84.97 84.29
-Swin-Unet [5] 86.56 88.06 83.07 82.41
-CMUNet [51] 88.33 89.48 84.82 85.56
-OWT 89.30 89.52 85.85 85.67
-OWT/${}_{\text{4slices}}$ 91.37 +2.03 91.41 +1.68 87.72 +1.87 87.31 +1.64
-Table 3: Performance of semantic segmentation.
+UNetplus 88.38 / 89.78 89.29 / 91.04 84.11 / 85.50 84.42 / 85.68
+AttU_Net 89.34 / 90.75 89.73 / 91.46 84.54 / 86.75 85.04 / 86.52
+TransUnet 88.91 / 90.24 89.70 / 91.34 84.97 / 86.40 84.29 / 86.01
+Swin-Unet 86.56 / 89.47 88.06 / 89.91 83.07 / 85.39 82.41 / 84.48
+CMUNet 88.33 / 90.92 89.48 / 91.50 84.82 / 86.82 85.56 / 86.94
+[…]
+Table 6: Dice performance (1- / 4-slice) of semantic segmentation.
+AutoPet Holistic Rec. 1 TG Rec. 1 TG Seg.
+[…]
+[…]
+AttU_Net 87.03
+CMUNet 87.89
+[…]
+[…]
+Table 7: Generalization capability of cross-dataset experiments trained on Abdomen1k and tested on AutoPet.
 Image Segmentation.
-We observe that reconstructing token groups with OWT preserves clear spatial details and boundary information in Fig.
-4.
-Consequently, a simple thresholding operation on the predictions can generate semantic binary masks, offering direct evidence of each organ-wise token group’s independence and interpretability.
+Reconstructing OWT’s token groups preserves clear spatial details and boundaries (Fig.
+4).
+Consequently, a simple thresholding operation on the predictions can generate semantic binary masks, offering direct evidence of each organ-wise token group’s independence and interpretability (More details on the thresholding operation are in Appendix B.3).
 Table 6 reports the single-organ segmentation Dice scores (averaged over the liver, kidneys, spleen, and pancreas).
-Overall, the accuracy of OWT’s organ-specific reconstructions is comparable to that of several dedicated 2D segmentation networks.
-Moreover, incorporating information from adjacent slices (OWT/${}_{\text{4slices}}$) further improves performance, surpassing the second-best approach by an average margin of 1.805 across four CT and MRI datasets (green number in Table 3).
-The detailed Dice of each organ are in Appendix 7.
-AutoPet Holistic-based Rec. 1 Token Group Rec. 1 Token Group Seg.
-L2 $\downarrow$ SSIM-3D $\uparrow$ L2 $\downarrow$ SSIM-3D $\uparrow$ Dice $\uparrow$
-MAE/${}_{\text{LPIPS}}$ [27] 3.58e-04 0.9872 - - -
-AttU_Net [42] - - - - 87.03
-CMUNet [51] - - - - 87.89
-OWT 1.12e-04 0.9909 4.07e-04 0.9765 87.54
-OWT/${}_{\text{4slices}}$ 9.34e-05 0.9924 3.60e-04 0.9798 89.42
-Table 4: Generalization capability of cross-dataset experiments trained on Abdomen1k and tested on AutoPet.
-Generalization.
-Table 4 reports the generalization performance when all methods are trained on Abdomen1k and tested on the AutoPet dataset.
-Under both reconstruction and segmentation metrics, OWT/${}_{\text{4slices}}$ outperforms dedicated holistic-based approaches.
-Moreover, it also achieves strong results in single-organ reconstruction, indicating that token groups can be flexibly combined or used individually to mitigate interference from irrelevant organs and background regions.
-Consequently, the semantic features extracted by OWT generalize more effectively, underscoring its potential as a foundational tokenization framework.
+For fair comparison, all networks report segmentation results based on 1 and 4 slices.
+Overall, the accuracy of OWT incorporating information from adjacent slices (OWT with 4-slice) achieves comparable or superior performance across CT and MRI datasets.
+Detailed per-organ Dice comparisons are in Appendix E.
+[…]
+4.4 Semantic Information in Token Groups
 Figure 6: t-SNE visualization of independent token groups on four medical datasets.
-4.2.2 Look Deeper into Token Groups
-Semantic Information in Token Groups.
+AutoPet Holistic Rec. 1 TG Rec.
+L2 $\downarrow$ LPIPS $\downarrow$ L2 $\downarrow$ LPIPS $\downarrow$
+OWT 1.49e-04 0.0232 3.32e-04 0.0071
+OWT w/ CT-CLIP 1.39e-04 0.0211 3.20e-04 0.0068
+Table 8: Reconstruction comparison between OWT w/ and w/o incorporated text prompts.
+Figure 7: Loss comparison between OWT w/ and w/o incorporated text prompts.
+T-SNE Analysis.
 Fig.
@@
 Within this space, token vectors of the same group type exhibit dense clusters, whereas different token groups remain distinctly separated.
-This clear inter-group separation highlights the semantic independence of organ-specific token groups in OWT, indicating strong potential for downstream tasks.
-Due to the semantic independence of the token group, Fig.
-5 shows that our approach not only supports holistic-level retrieval, but also enables organ-level retrieval ($L_{2}$ distance measure).
-Specifically, we demonstrate top-1, top-2, and maximum-distance retrieval results on both CT and MRI modalities for different token groups.
-The case indices in the top-left corner of each row indicate how retrieval results vary for different organs of interest.
-Since experiments in reconstruction and segmentation demonstrate that each token group preserves both the spatial structure and texture details of its corresponding organ, this organ-specific retrieval can be highly beneficial in medical scenarios.
-For example, it can help radiologists quickly locate similar organ regions across multiple patients, facilitate region-based comparisons of pathology progression, or assist in reference image selection for organ-specific diagnoses.
-AutoPet Holistic-based Rec. 1 Token Group Rec.
+This clear inter-group separation highlights the semantic independence of organ-specific token groups in OWT.
+Multi-Modality Incorporation.
+OWT’s highly disentangled and independent SDTGs enable efficient alignment or incorporation with other modalities, such as text.
+We demonstrate this by encoding medical prompts with a CT-CLIP encoder (specifically pretrained on paired CT images and reports) [hamamci2024developing], and adding the resulting organ-specific and CT-consistent text embeddings element-wise to their corresponding token groups.
+More details on prompts and architecture for training OWT with text encoding are in Appendix C.
+As shown in Fig.
+7 and Table 8, this simple integration accelerates loss convergence and provides immediate improvements in several downstream tasks.
+This highlights OWT’s promise for multi-modal enhancement, unlocking potential applications, such as text-guided retrieval, generation, and improved interpretability in complex workflows.
+AutoPet Data Holistic Rec. 1 TG Rec. Seg.
+[…]
+OWT 100% 1.49e-04 0.0232 3.32e-04 0.0071 0.8952
+OWT 20% 5.54e-04 0.0658 5.70e-04 0.0115 0.8109
+[…]
+OWT 60% 1.86e-04 0.0270 3.49e-04 0.0076 0.8858
+[…]
+Table 9: Performance comparison of OWT in different training strategies.
+CelebAMaskHQ Holistic Rec. 1 Token Group Rec.
 L2 $\downarrow$ LPIPS $\downarrow$ L2 $\downarrow$ LPIPS $\downarrow$
-OWT 1.49e-04 0.0232 3.32e-04 0.0071
-OWT w/ Text 1.41e-04 0.0221 3.16e-04 0.0068
-Table 5: Reconstruction comparison between OWT w/ and w/o incorporated text prompts.
-Figure 7: Loss comparison between OWT w/ and w/o incorporated text prompts.
-Multi-Modality Incorporation.
-In OWT, each organ is embedded into its corresponding SDTG, resulting in highly semantically disentangled and independent representations.
-Therefore, these representations can be efficiently aligned with or collaborated on with other modality embeddings (e.g., text encoding).
-Specifically, during OWT training, we use medical prompts and a pretrained CLIP encoder [45] to obtain text encoding of each organ, which are then added element-wise to their matching token groups after the Organ Collector.
-(More details on prompts and architecture for training OWT w/ text encoding are in Appendix 6.2.2.)
-Fig.
-7 and Table 5 reveal that simply adding text prompts to organ-wise token groups instantly accelerates loss convergence and provides slight improvements in both holistic- and semantic-based reconstruction.
-Consequently, OWT shows promise for incorporating the text modality to further enhance organ-specific representations.
-Such multi-modal integration could unlock new medical applications, such as more robust organ labeling, guided retrieval and generation, and improved interpretability in complex workflows.
-4.2.3 Organ-wise Token Groups are more than Organs
-CelebAMaskHQ Holistic-based Rec. 1 Token Group Rec.
-L2 $\downarrow$ LPIPS $\downarrow$ L2 $\downarrow$ LPIPS $\downarrow$
-MAE/${}_{\text{LPIPS}}$ [27] 2.78e-03 0.1157 - -
-OWT 2.50e-03 0.1101 2.83e-03 0.0458
+[…]
+[…]
 Table 10: Performance of OWT on CelebAMaskHQ dataset.
+Figure 8: Samples of holistic and semantic-based human face reconstruction on CelebAMaskHQ.
+4.5 Semi-Supervised Training under Limited Annotations
+The limitation of annotation is a long-standing issue in medical imaging.
+In our foundational framework, we find that appropriate supervision of token groups is beneficial for downstream tasks.
+Therefore, we adopted a two-step training strategy: (1) using all data without labels for unsupervised pretraining of OWT’s encoder and decoder via whole image reconstruction for 600 epochs, and (2) fine-tuning the entire OWT with only partial annotation coverage (20% and 60%) from the training set for 1200 epochs.
+As shown in Table 9, this approach significantly improves performance under label scarcity.
+The semi-supervised model using 20% labels greatly outperformed the model trained on only 20% labels.
+Furthermore, the semi-supervised mode with 60% labels nearly matched the upper-bound results of training with 100% labels , showing OWT’s robustness and potential for data-scarce scenarios.
+4.6 Token Group as Compositional Units
+Figure 9: Semantic-based retrieval of CelebAMaskHQ.
 Although OWT and TGR are primarily designed for grouping different semantic components in medical images, they can also be applied to non-medical datasets.
@@
 8 and Table 10 that OWT successfully learns semantically disentangled token groups on the CelebAMaskHQ dataset, capturing and reconstructing each facial region with high spatial and textural fidelity.
-Figure 8: Samples of holistic-based and semantic-based human face reconstruction on CelebAMaskHQ.
-Figure 9: Semantic-based retrieval of CelebAMaskHQ.
 As illustrated in Fig.
@@
 Our approach not only improves interpretability, generalizability, efficiency and controllability in organ-based analyses, but also shows adaptability to broader scenarios, including natural image domains.
-By enabling semantic-level token groups, OWT enriches downstream tasks such as semantic-based reconstruction, segmentation and retrieval, offering a new perspective on how representations can be leveraged.
+By enabling semantic-level token groups, OWT enriches unique clinical applications and downstream tasks such as organ-specific analysis, semantic-based reconstruction, segmentation and retrieval, offering a new perspective on how representations can be utilized.
+[…]
 Supplementary Material
-Abdomen1k Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
-UNetplus [63] 96.85 94.63 91.00 71.05 88.38
-AttU_Net [42] 97.09 95.22 93.05 72.00 89.34
-TransUnet [9] 96.78 94.65 92.30 71.89 88.91
-SwinUnet [5] 96.53 93.68 90.52 65.50 86.56
-CMUNet [51] 96.05 94.67 90.84 71.75 88.33
-[…]
-OWT/${}_{\text{4slices}}$ 96.81 94.84 95.59 78.23 91.37 +2.03
-AutoPet Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
-UNetplus [63] 96.41 92.56 91.31 76.87 89.29
-AttU_Net [42] 96.60 92.86 91.22 78.24 89.73
-TransUnet [9] 96.34 93.59 90.65 78.23 89.70
-SwinUnet [5] 96.33 92.34 91.11 72.46 88.06
-CMUNet [51] 96.56 93.11 90.61 77.65 89.48
-[…]
-OWT/${}_{\text{4slices}}$ 96.65 93.15 93.96 81.87 91.41 +1.68
-Table 7: Performance of semantic segmentation on CT datasets.
-Delay Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
-UNetplus [63] 93.53 91.85 85.20 65.85 84.11
-AttU_Net [42] 94.25 91.08 86.85 65.98 84.54
-TransUnet [9] 93.93 90.83 89.07 66.03 84.97
-SwinUnet [5] 92.78 90.61 86.35 62.55 83.07
-CMUNet [51] 94.02 91.88 87.74 65.64 84.82
-[…]
-OWT/${}_{\text{4slices}}$ 94.72 92.61 92.23 71.33 87.72 +1.87
-PreArtery Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
-UNetplus [63] 94.05 92.25 87.26 64.14 84.42
-AttU_Net [42] 94.33 92.32 89.07 64.45 85.04
-TransUnet [9] 94.33 90.69 86.59 65.56 84.29
-SwinUnet [5] 92.91 88.91 86.33 61.48 82.41
-CMUNet [51] 94.26 92.58 89.21 66.18 85.56
-[…]
-OWT/${}_{\text{4slices}}$ 94.74 92.18 92.14 70.17 87.31 +1.64
-Table 8: Performance of semantic segmentation on MRI datasets.
+Appendix A Related Work
+A.1 Holistic Embedding-based Representation Learning
+Holistic embedding-based approaches map entire 2D/3D inputs into a single latent representation, aiming to capture global image features [voulodimos2018deep, geirhos2020shortcut, wang2024disentangled].
+These methods typically follow either self-supervised or task-guided strategies.
+In self-supervised learning, contrastive techniques like SimCLR [chen2020simple], SimSiam [chen2021exploring], and BYOL [grill2020bootstrap] enforce consistency between augmented views, while predictive approaches such as CPC [oord2018representation] and RotNet [gidaris2018unsupervised] reconstruct missing or permuted portions.
+Recently, masked image modeling (e.g., MAE [he2022masked] and BEiT [bao2021beit]) has emerged as an effective approach by requiring models to infer masked regions.
+Meanwhile, task-guided frameworks focus on extracting features that directly benefit a target task, typically resulting in holistic embeddings.
+Many classic neural networks, such as ResNet [he2016deep], DenseNet [huang2017densely], and Vision Transformer [dosovitskiy2020image], and their variants incorporate classification, regression, or segmentation objectives.
+In medical imaging, holistic representations have been widely used for self-supervised feature learning in CT [chen2021momentum, lai2024e3d], MRI [chaitanya2020contrastive, lang20233d], and X-ray [tiu2022expert, xiao2023delving].
+Additionally, classical supervised pipelines (e.g., U-Net [ronneberger2015u] and its variants [zhou2018unet++, chen2021transunet, cao2022swin, hatamizadeh2021swin]) rely on a global bottleneck to embed entire images or volumes, demonstrating strong performance in organ segmentation and disease detection.
+Although such approaches achieve notable success, they often encode multiple semantic factors, such as multiple organs and the surrounding background, into an entangled representation with limited interpretability and fine-grained control [schutte2021using, cetin2023attri, wang2024disentangled].
+A.2 Disentangled Representation Learning
+In contrast to holistic embedding-based approaches, various disentangled representation learning (DRL) studies have pursued a dimension-wise strategy for two main objectives.
+First, to derive more robust holistic embeddings with improved generalizability [dangovski2021equivariant, bai2023robust, devillers2023equimod, wang2024distortion, garrido2023self, song2024contrastive, wang2024disentangled], and second, to explicitly decompose images into latent factors, each corresponding to distinct and semantically meaningful attributes, in order to improve explainability [burgess2018understanding, chen2016infogan, xiao2017dna, zhu2018visual].
+In medical imaging, the motivations for DRL are especially strong: improved interpretability, controllability, and robustness to confounding factors are crucial in clinical settings [liu2022learning, wang2022disentangled].
+When transitioning DRL to medical imaging, researchers seek to decouple anatomical properties (e.g., organ size or shape) [wang2022semantic, jiang2022disentangled], modality characteristics (e.g., CT and MRI sequences) [ouyang2021representation, chartsias2019disentangled], and pathological factors (e.g., tumor presence or severity) [liu2022joint, hu2022clinically], aiming to isolate distinct clinical features.
+Despite notable progress, data- and perturbation-driven DRL approaches can be sensitive to training objectives, architecture choices, and assumptions about independent factors [locatello2019challenging, wang2024distortion, wang2024disentangled].
+In existing studies, disentangled dimensions may not align precisely with real anatomical or pathological concepts, and spatial correspondence often remains weak when images are globally encoded.
+Consequently, there has been a growing emphasis on methods that combine disentanglement principles with controlled grouping or tokenization schemes to encapsulate semantically meaningful subsets of features.
 Figure 10: Comparison of segmentation performance using direct and indirect approaches.
-6 Implementation Details
-6.1 Image Preprocessing
-For all 3D medical volumes, we standardize the spatial resolution to $2\text{mm}\times 2\text{mm}\times 2\text{mm}$ and apply a center crop of $224\times 224$.
-For CT datasets, we clip intensity values to the $[-1000,1000]$ Hounsfield Unit (HU) range and normalize them to $[0,1]$.
+Appendix B Implementation Details
+B.1 Image Preprocessing
+For all 3D medical volumes, we standardize the spatial resolution to $2.0\text{mm}\times 2.0\text{mm}\times 2.0\text{mm}$ and apply a center crop of $224\times 224$.
+For CT datasets, we clip intensity values to the [1000,1000][-1000,1000] Hounsfield Unit (HU) range and normalize them to $[0,1]$.
 For MRI datasets, we clip intensities to the $[0,99]$ percentile range and normalize them to $[0,1]$.
@@
 Each medical image is further standardized in thickness by center-padding or cropping 112 slices based on the anatomical centers of the four organs.
-For the Abdomen1K dataset [40], the number of cases in the training and test sets is 899:101.
-For AutoPet [21], the split is 700:200, while for RAOS delay and pre-artery sequences [39], both have a training-to-test ratio of 330:83.
+For the Abdomen1K dataset [ma2021abdomenct], the number of cases in the training and test sets is 899:101.
+For AutoPet [Gatidis2022], the split is 700:200, while for RAOS delay and pre-artery sequences [luo2024rethinking], both have a training-to-test ratio of 330:83.
 All experiments, including OWT, MAE, and segmentation networks, are based on the same preprocessed dataset.
-For the 2D face images in CelebAMaskHQ dataset [34], we resize them to $224\times 224$, and define six semantic entities based on annotations (background, face&ear, hair&hat, mouth, nose, and eyes&glasses), as shown in Fig.
+For the 2D face images in CelebAMaskHQ dataset [lee2020maskgan], we resize them to $224\times 224$, and define six semantic entities based on annotations (background, face&ear, hair&hat, mouth, nose, and eyes&glasses), as shown in Fig.
 8.
 The CelebAMaskHQ dataset contains 30,000 facial image cases, with 24,000 images allocated for training and 6,000 for testing.
-6.2 Architecture Details
-6.2.1 Default Architecture of MAE and OWT
-In OWT, we adopt the same encoder and decoder architecture as MAE, with the only modification being the replacement of the self-attention blocks [53] with linear self-attention blocks [31].
+B.2 Default Architecture of OWT and MAE
+In OWT, we adopt the same encoder and decoder architecture as MAE [he2022masked], with the only modification being the replacement of the self-attention blocks [vaswani2017attention] with linear self-attention blocks [katharopoulos2020transformers].
+The hyperparameters, optimizer and scheduler settings all follow MAE.
 By default, we use the ViT-base architecture, where the MAE encoder consists of 12 attention blocks, the decoder has 8, and the token embedding dimension is set to 768.
@@
 This ensures that the overall model length remains consistent with MAE.
-Additionally, since each SDTG contains a default of 20 tokens, the computational load of OWT is lower than that of MAE when using fewer than 10 SDTGs.
+The default token number for each token group is 20 in OWT.
+Therefore, the training load of OWT is lower than that of MAE when using fewer than 10 SDTGs.
 To prevent information leakage during TGR training, we remove the class token from the OWT architecture.
@@
 For 3D input volumes, we apply a $1\times 16\times 16$ 3D kernel for patch embedding.
-6.2.2 OWT w/ Text Encoding
-For the OWT experiment incorporating text encoding, we design a text prompt inspired by [35].
-Specifically, we define the template as "a computerized tomography [category]", where [category] corresponds to "background without the Liver, Kidneys, Spleen, and Pancreas", "of a Liver", "of Kidneys", "of a Spleen", and "of a Pancreas", aligning with the number of token groups.
-These prompts are fed into a pretrained CLIP text encoder [45] to obtain the corresponding text encoding, resulting in a representation of size $5\times 512$.
-To match the token groups, we replicate each text encoding according to the number of tokens per token group (20), leading to a final text encoding dimension of $100\times 512$.
-Within the OWT framework, a linear transformation layer maps the text encoding from $512$ to $768$, producing a transformed text embedding of size $100\times 768$.
-This transformed text encoding is then element-wise added to the organ-wise representation $X_{G}$ extracted by the Organ Collector.
-The combined representation proceeds through the random selection procedure in TGR, while all other components remain consistent with the default OWT settings.
-6.3 Experimental Setup
+Method AutoPet CelebAMaskHQ
+L2 $\downarrow$ LPIPS $\downarrow$ L2 $\downarrow$ LPIPS $\downarrow$
 […]
+OWT 1.49e-04 0.0232 2.50e-03 0.1101
+[…]
+Table 11: Performance comparison when altering the Transformer-based encoder and decoder to those of VQGAN.
+B.3 Experimental Setup
+We follow the same experimental setup and hyperparameter settings as MAE.
+Additionally, each MAE- and OWT-based experiment is trained for 1200 epochs, with a 60-epoch warm-up phase.
+The training follows the learning rate protocol $\text{lr}=\text{base\_lr}\times{\text{effective\_batch\_size}}/{256}$ where the base learning rate is set to $\text{base\_lr}=10^{-4}$.
 For OWT with 2D image inputs, we set the batch size to 96.
@@
 For the interpretability experiments validated by segmentation, we apply a 0.15 threshold to transfer the reconstructed results as semantic masks.
-7 Detailed Semantic Segmentation Results
-Detailed Results.
-Tables 7 and 8 provide a more detailed results of those presented in Table 3, showing the Dice scores for each of the four organs across different datasets.
-From Tables 7 and 8, we observe that compared to dedicated segmentation networks, OWT/${}_{\text{4slices}}$ generally achieves superior performance across all organs.
-Notably, for smaller organs such as the pancreas, OWT/${}_{\text{4slices}}$ demonstrates a significant improvement, with a Dice score advantage exceeding 5.3 in the best case.
+Appendix C Detailed Settings of Multi-Modality Incorporation
+For the OWT experiment incorporating text encoding, we design a text prompt inspired by [liu2023clip].
+Specifically, we define the template as "a computerized tomography [category]", where [category] corresponds to "background without the Liver, Kidneys, Spleen, and Pancreas", "of a Liver", "of Kidneys", "of a Spleen", and "of a Pancreas", aligning with the number of token groups.
+These 5 prompts are fed into a CT-specific pretrained CLIP text encoder, CT-CLIP [hamamci2024developing] to obtain the corresponding CT-consistent text encoding, resulting in a representation of size $5\times 512\times 768$.
+Then, the mean is taken along the middle dimension, yielding a representation of size $5\times 768$.
+To match the dimension token groups, we replicate each text encoding according to the number of tokens per token group (20), leading to a final text encoding dimension of $100\times 768$.
+Within the OWT framework, a linear transformation layer maps the text encoding from $768$ to $768$, resulting in a transformed text embedding of the same size $100\times 768$.
+This transformed text encoding is then element-wise added to the organ-wise representation $X_{G}$ extracted by the Organ Collector.
+The combined representation proceeds through the random selection procedure in TGR, while all other components remain consistent with the default OWT settings.
+Appendix D Ablation Studies
+For the ablation study, we explore the following key aspects: the flexibility of the encoder-decoder architecture, the number of tokens per SDTG, and the impact of varying the amount of input data and labels in OWT.
+D.1 Backbone Alteration
+We investigate the effect of replacing the default ViT-based encoder-decoder architecture in OWT (adopted from MAE) with the encoder and decoder from VQGAN [esser2021taming].
+Specifically, we configure the VQGAN bottleneck dimensions to $14\times 14$ for a $224\times 224$ input image, ensuring that the number of latent tokens (196) remains consistent with the ViT-based setting.
+As shown in Table 11, despite the increased computational cost of VQGAN’s hybrid CNN and self-attention structure, it achieves significant performance gains over the transformer-based encoder-decoder.
+On two datasets, it improves both L2 and LPIPS evaluation metrics by 53.8%–68.5% in AutoPet and 43.8%–51.0% in CelebAMaskHQ, demonstrating the potential for further enhancements in OWT and highlighting its architectural flexibility.
+Figure 11: Ablation study on the number of tokens in each token group when trained on AutoPet.
+AutoPet Holistic Rec. 1 Token Group Rec.
+OWT L2 $\downarrow$ LPIPS $\downarrow$ L2 $\downarrow$ LPIPS $\downarrow$
+Regular 1.49e-04 0.0232 3.32e-04 0.0071
+Adaptive 1.40e-04 0.0215 3.19e-04 0.0067
+Table 12: Reconstruction comparison of OWT with regular and adaptively distributed SDTGs.
+D.2 Number of Tokens per SDTG
+The number of tokens within each SDTG determines its capacity to store organ-specific information.
+To evaluate this, we experiment with SDTG token counts of 5, 10, 20, and 40, assessing their impact on both generation and segmentation tasks.
+For the generation task, we measure full-image reconstruction using all SDTGs and single-organ reconstruction using only one SDTG, evaluated by L2, LPIPS, and SSIM-3D metrics.
+For segmentation, we assess the Dice score of single-organ masks.
+As shown in Fig.
+11, increasing the token count from 10 to 20 provides a slight performance improvement, while further increasing it from 20 to 40 leads to a more significant boost.
+This suggests that OWT has considerable scalability and potential as a foundational semantic-level tokenization framework, capable of further enhancements with increased token capacity.
+Figure 12: Ablation study on OWT performance when utilizing different percentages of the dataset (AutoPet).
+Abdomen1k Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
+UNetplus 96.85 94.63 91.00 71.05 88.38
+AttU_Net 97.09 95.22 93.05 72.00 89.34
+TransUnet 96.78 94.65 92.30 71.89 88.91
+SwinUnet 96.53 93.68 90.52 65.50 86.56
+CMUNet 96.05 94.67 90.84 71.75 88.33
+OWT 96.55 94.21 95.12 71.32 89.30
+[…]
+AutoPet Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
+UNetplus 96.41 92.56 91.31 76.87 89.29
+AttU_Net 96.60 92.86 91.22 78.24 89.73
+TransUnet 96.34 93.59 90.65 78.23 89.70
+SwinUnet 96.33 92.34 91.11 72.46 88.06
+CMUNet 96.56 93.11 90.61 77.65 89.48
+OWT 96.14 92.33 92.79 76.82 89.52
+[…]
+Table 13: Performance of semantic segmentation on CT datasets.
+Delay Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
+UNetplus 93.53 91.85 85.20 65.85 84.11
+AttU_Net 94.25 91.08 86.85 65.98 84.54
+TransUnet 93.93 90.83 89.07 66.03 84.97
+SwinUnet 92.78 90.61 86.35 62.55 83.07
+CMUNet 94.02 91.88 87.74 65.64 84.82
+OWT 94.32 91.87 91.55 65.65 85.85
+[…]
+PreArtery Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
+UNetplus 94.05 92.25 87.26 64.14 84.42
+AttU_Net 94.33 92.32 89.07 64.45 85.04
+TransUnet 94.33 90.69 86.59 65.56 84.29
+SwinUnet 92.91 88.91 86.33 61.48 82.41
+CMUNet 94.26 92.58 89.21 66.18 85.56
+OWT 94.19 91.10 91.82 65.57 85.67
+[…]
+Table 14: Performance of semantic segmentation on MRI datasets.
+D.3 Adaptively Distributed SDTGs
+To explore this, we include a proof-of-concept experiment showing that the OWT framework has the potential to support adaptive token group sizing.
+We allocated token numbers to different organs based on the fourth root of their volumes.
+For background (containing all information other than defined organs), liver, kidney, spleen, and pancreas, 45, 20, 13, 13, and 9 were allocated, with the total tokens fixed at 100 (identical to the default setting).
+Preliminary results show that this adaptive allocation better accommodates the varying feature distributions across organs and provides performance improvements over a fixed-size setting (Table 12), demonstrating the scalability of OWT as a novel foundational tokenization framework.
+D.4 Data Percentage
+Fig.
+12 illustrates the performance of the OWT framework on the AutoPet dataset when trained with 20%, 60%, and 100% of the data.
+We observe a significant improvement in performance when increasing the dataset size from 20% to 60%, whereas the gain from 60% to 100% is relatively marginal.
+This suggests that OWT efficiently learns meaningful representations with a moderate amount of data, indicating its strong data efficiency and the potential for effective generalization even with limited training samples.
+Figure 13: Semantic-based retrieval of CT images.
+Appendix E Detailed Semantic Segmentation Results
+[…]
 Token Group Connectivity.
@@
 6, further validating their disentangled nature.
-Method AutoPet CelebAMaskHQ
-L2 $\downarrow$ LPIPS $\downarrow$ L2 $\downarrow$ LPIPS $\downarrow$
-MAE/${}_{\text{LPIPS}}$ 2.77e-04 0.0255 2.78e-03 0.1157
-OWT 1.49e-04 0.0232 2.50e-03 0.1101
-OWT/${}_{\text{VQGAN}}$ 6.89e-05 ↓53.8% 0.0073 ↓68.5% 1.80e-03 ↓43.8% 0.0540 ↓51.0%
-Table 9: Performance comparison when altering the Transformer-based encoder and decoder to those of VQGAN.
-8 Ablation Studies
-For the ablation study, we explore the following key aspects: the flexibility of the encoder-decoder architecture, the number of tokens per SDTG, and the impact of varying the amount of input data in OWT.
-Backbone Alteration.
-We investigate the effect of replacing the default ViT-based encoder-decoder architecture in OWT (adopted from MAE) with the encoder and decoder from VQGAN.
-Specifically, we configure the VQGAN bottleneck dimensions to $14\times 14$ for a $224\times 224$ input image, ensuring that the number of latent tokens (196) remains consistent with the ViT-based setting.
-As shown in Table 9, despite the increased computational cost of VQGAN’s hybrid CNN and self-attention structure, it achieves significant performance gains over the transformer-based encoder-decoder.
-On two (medical and facial) datasets, it improves both L2 and LPIPS evaluation metrics by 43.8%–68.5%, demonstrating the potential for further enhancements in OWT and highlighting its architectural flexibility.
-Figure 11: Ablation study on the number of tokens in each token group when trained on AutoPet.
-Figure 12: Ablation study on OWT performance when utilizing different percentages of the dataset (AutoPet).
-Number of Tokens per SDTG.
-The number of tokens within each SDTG determines its capacity to store organ-specific information.
-To evaluate this, we experiment with SDTG token counts of 5, 10, 20, and 40, assessing their impact on both generation and segmentation tasks.
-For the generation task, we measure full-image reconstruction using all SDTGs and single-organ reconstruction using only one SDTG, evaluated by L2, LPIPS, and SSIM-3D metrics.
-For segmentation, we assess the Dice score of single-organ masks.
-As shown in Fig.
-11, increasing the token count from 10 to 20 provides a slight performance improvement, while further increasing it from 20 to 40 leads to a more significant boost.
-This suggests that OWT has considerable scalability and potential as a foundational semantic-level tokenization framework, capable of further enhancements with increased token capacity.
-Data Percentage.
-Fig.
-12 illustrates the performance of the OWT framework on the AutoPet dataset when trained with 20%, 60%, and 100% of the data.
-We observe a significant improvement in performance when increasing the dataset size from 20% to 60%, whereas the gain from 60% to 100% is relatively marginal.
-This suggests that OWT efficiently learns meaningful representations with a moderate amount of data, indicating its strong data efficiency and the potential for effective generalization even with limited training samples.
+Appendix F Additional Organ-Level Retrieval Demonstration
+An additional example demonstrates OWT performing organ-level retrieval on the CT modality (Fig.
+13).
+Combined with Fig.
+5, this shows the stability and potential of OWT for organ-level retrieval across different modalities.
+It can not only retrieve patients with similar body shapes at a holistic level, but also retrieve organs in similar conditions and locations for specific organs of interest.
+This highlights the more flexible and unique clinical value of OWT, which cannot be achieved by existing retrieval methods based on holistic embeddings.